Linux Commercial-HOWTO

Martin Michlmayr, tbm@sypher.com, http://www.sypher.com/tbm v2.0, 29 June 1996 This document contains a listing of commercial software and applications which are offered for Linux. It is maintained by Martin Michlmayr <tbm@sypher.com>.

Contents

1 About this document 5

2 Copyright Information 6

3 Disclaimer 7

4 Related Information 7

5 Marketing aspects of commercial Linux applications 7 5.1 Investigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.2 Some Aspects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6 Linux for the Commercial Market? 11 6.1 Turnkey systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.2 Large end-user customers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.3 Specialized markets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.4 Biography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

7 USELINUX Conference 16 7.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.2 Tutorials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.3 Technical track topics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.4 Business track topics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.5 How to submit a proposal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 7.6 Preferred method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 7.7 Alternative method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 7.8 Vendor exhibits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.9 Birds-Of-a-Feather sessions (BOFs) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.10 About the USENIX Association . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.11 About Linux International . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.12 Registration information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

8 Linux vs. other free Unix flavors 20

9 Product Groups 21

9.1 Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 9.1.1 AccountFlex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 9.1.2 CONZEPT 16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 9.1.3 D-ISAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 9.1.4 Empress RDBMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 9.1.5 ESQLFlex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 9.1.6 InfoFlex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.1.7 Just Logic/SQL Database Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 9.1.8 POET 2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 9.1.9 /rdb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.1.10 Veritas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.1.11 Yard SQL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

9.2 Data Visualisation, CAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 9.2.1 Ghostscript 3.x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 9.2.2 IDL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 9.2.3 MRJ Symbolic OCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 9.2.4 SISCAD-P 1.3-3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 9.2.5 TecPlot 6.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 9.2.6 XVScan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

9.3 Development tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 9.3.1 Amzi! Prolog & Logic Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 9.3.2 Basmark QuickBASIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 9.3.3 Clickables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 9.3.4 CODINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 9.3.5 Distributed Interface Object Server System (DIOSS) . . . . . . . . . . . . . . . . . . . 47 9.3.6 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 9.3.7 EZ Clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

9.3.8 Dynace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

9.3.9 Eiffel-S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

9.3.10 Finesse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

9.3.11 Flagship . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

9.3.12 Gen/X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

9.3.13 ICC11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

9.3.14 INSURE++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

9.3.15 INTERACTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

9.3.16 ISE Eiffel 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

9.3.17 Metacard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

9.3.18 Mjolner BETA System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

9.3.19 SEDIT, S/REXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

9.3.20 Smalltalk/X language system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

9.3.21 tgdb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

9.3.22 TowerEiffel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

9.4 Financial Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

9.4.1 BB Tool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

9.5 Mathematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

9.5.1 Maple V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

9.5.2 Mathematica . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

9.5.3 Matlab and Simulink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

9.5.4 Reduce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

9.6 Network Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

9.6.1 Galacticomm BBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

9.6.2 NetEye . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

9.6.3 Venus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

9.6.4 VU-BBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

9.7 Text Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

9.7.1 CRiSP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

9.7.2 ibgsXaed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

9.8 X Windows related products . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 9.8.1 BXwidgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 9.8.2 BXwidgets/DB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 9.8.3 Metrolink Motif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.8.4 MoCheap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.8.5 OpenGL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 9.8.6 X Inside X Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

9.9 Other Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 9.9.1 Abuse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 9.9.2 Amadeus music software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 9.9.3 aqua zis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 9.9.4 ChemApp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 9.9.5 CLiX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 9.9.6 Emulus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 9.9.7 Executor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 9.9.8 Mazama Packet Filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 9.9.9 PROCHEM-C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 9.9.10 ROSIN NC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 9.9.11 SPATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 9.9.12 Unix Cockpit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 9.9.13 Unix Auto-Tutor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

9.10 Free Software for Commercial Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 9.10.1 Cyclades Cyclom-Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 9.10.2 Stallion Technologies Multiport Serial Boards . . . . . . . . . . . . . . . . . . . . . . . 111

9.11 German corner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 9.11.1 BBSD-Druckerei . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 9.11.2 Orgaplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

10 Support 116 10.1 Companies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

1 About this document

This is the Linux Commercial-HOWTO. It contains a listing of commercial software which is available for Linux. The Commercial-HOWTO doesn’t contain any information on Linux distributions – this is covered by the Distribution-HOWTO.

This document was originated by Harald Milz <Harald.Milz@Linux.org>. It is now maintained by Martin Michlmayr <tbm@sypher.com>.

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>.

Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Commercial-HOWTO is available via FTP at ftp://ftp.sypher.com/tbm/HOWTO or can be found on the World Wide Web at http://www.sypher.com/tbm/HOWTO.

The Commercial-HOWTO is not a forum for product announcements or marketing hype; it is a service for potential customers and the whole Linux community. Resellers will not be listed; the list is for companies who produce their software themselves. Two main goals are being aimed at:

• It shall help companies who want to run Linux to find software solutions and applications. The international distribution of this list will enhance the contact opportunity.

• It is meant to prove the commercial useability of Linux and thus to encourage other vendors to port their software as well.

Companies and developers who are offering their products for Linux and interested in joining the Commercial-HOWTO are invited to fill in the following form and contact me at tbm@sypher.com.

This HOWTO contains tabular entries for each product (example follows). The entry format is similar to the Linux Software Map (LSM) entry (field/stanza lengths are arbitrary). If you want me to add your entry please keep short, otherwise I’ll have to shorten your data. Furthermore, please send me plain ASCII data; no HTML, no Postscript, and no MIME.

If you send me an entry for your product, please don’t forget to send me timely updates if appropriate.

     Name:
     The name of your application.
     Description:
     SHORT description of the package, just the basic functionality.
     Licensing Policy:
     Whatever applies.  Is there a free demo or shareware version available
     via FTP or WWW?  Where?

     OS Provisions:
     Kernel version, XFree86 version, Motif version, RAM, harddisk usage, etc.
     Documentation:
     Printed documentation, page number, online help, language.
     Extra Features and Add-Ons:
     (and their prices)
     Price Range:
     Whatever it costs.
     Vendor:
     Address:
     Phone: (U.S. and Canada: if you like to be reachable internationally,
            please don't enter only a 1-800 or 1-900 number)
     Fax:
     Contact:
     EMail:
     WWW:

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 by Martin Michlmayr. A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it. Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Commercial-HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions. In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs. We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 Disclaimer

This HOWTO is not actually a HOWTO in the sense of the Linux Documentation Project. Instead, it is an instrument to investigate the commercial Linux opportunity and to list applications which were already ported and marketed in a native Linux version. As a software vendor, you probably know that you can alternatively offer Linux users a statically linked SCO version of your application which would probably run under the iBCS2 emulator (albeit with a small performance penalty and higher memory requirements). Such applications will not be listed here (see the iBCS2 documentation at ftp://tsx-11.mit.edu/pub/linux/BETA/ibcs2).

I will not select nor deselect any particular product. Instead, everyone who wants to have her product included will be serviced. However, I reserve the right to shorten individual entries to keep things in shape.

If you don’t find a particular product or vendor in this list, this is probably due to one of the following reasons:

• I never heard of that product or vendor and thus didn’t try to get in contact.

• I did get in contact, but the vendor didn’t answer yet.

• I did get in contact, but the vendor stated positively that he doesn’t sell his product for Linux (yet).

In any case, please get in contact if you feel someone’s missing; also if you discover any errors in the file.

Sometimes two vendor’s addresses are mentioned in the “Vendor:” field. In these cases I received the information from the German subsidiary/distributor. The original manufacturer’s address is always mentioned first.

My comments in the text are enclosed in parentheses () and marked with my initials “tbm”.

4 Related Information

There’s another HTML page which covers commercial Linux software. It is maintained by Alan Cox <Alan.Cox@linux.org> and can be found at http://www.linux.org.uk/LxCommercial.html.

5 Marketing aspects of commercial Linux applications

This section was suggested by Orest Zborowski <orestz@eskimo.com>; it is still under construction.

It is meant to help potential commercial software developers to market their products. There are several caveats to watch when doing this, e.g. the GNU General Public License (GPL) and the GNU Library General Public License (GLPL).

I’d appreciate anyone who wants to share her experience with Linux-specific marketing issues to send me some lines (see questionaire below).

Please make sure that people interested in more specific items can contact you via EMail or fax.

5.1 Investigation

Here is a questionaire which I ask you to fill out and send me. Answers will be handled anonymously, and I will only disclose facts, no names. The facts you tell me will be of great value to get an idea how the commercial Linux market looks like. If you do not market your application in a native Linux version, please tell my why. If not stated otherwise, “application” stands for “native Linux version of your application”.

• Why did you start marketing your application? Why didn’t you do it earlier?

• How many copies of the application do you expect to sell – for Linux,

– for all the other Unix variants, during the planned product life? (This information is kept confidential if you require)

• How many copies of your application did you already sell – for Linux?

– for all the other Unix variants? (This information is kept confidential if you require)

• Which technical problems did you have porting your application to Linux? How did you solve/circumvent them?

• How many employees are working on – porting to Linux,

– providing support to end users/distributors? (This information is kept confidential if you require)

• Does your company offer the support for the application or did you (attempt to) outsource the support? If you provide support yourself, would you like to outsource it?

• Which issues were important for the pricing? Is the application – more expensive,

– equally priced,

– cheaper,

– shareware priced, compared to versions for other Unix variants?

• What types of end users you you address with the application?

• According to your experience, was it a good decision to start marketing the application? Will you port and market other applications too?

• Are there any copy protection issues you have/like to consider? If yes, how did you solve the problem?

• How do you think about the iBCS2 emulator and offering a statically linked SCO/ISC or other version runnable under the emulator? Why did you choose to offer a native Linux version instead?

• What is your warranty strategy (Do you restict the usage to certain kernel/C-lib versions etc.)?

• Which country-specific (e.g. legal) issues did you have to take into consideration? How did you cope with them?

• Are there any other important issues which I possibly (probably) forgot to ask?

5.2 Some Aspects

Porting to Linux simply means recompiling in many, many cases. Sometimes some #include tweaking may be necessary. If you develop your apps using GNU C/C++ which is very useful sometimes, there should be virtually no further work to do.

One big advantage with Linux is that all the source code of the operating system is freely available. Experienced programmers can easily spot the locations where problems occur and contact the ”responsible” developer for assistance. I never encountered anyone being impolite – just make sure you read the respective documentation first...

The persons responsible for the Linux port should know where to get vital information in case of problems. The Linux Documentation Project offers plenty information about all and everything and the kitchen sink. The Linux INFO-SHEET is a good place to start with for beginners.

The key to success with your Linux port is Internet access, particularly FTP access to the big Linux FTP servers such as sunsite.unc.edu or tsx-11.mit.edu and their legion mirrors.

It appears that one big problem is after-sales support. If you do market your software for Linux, you should consider to employ someone who knows whom and where to ask in case of problems with the OS itself. On the other hand, you may expect problems with Linux to be solved much faster than problems with commercial Unix variants. For example, the /bin/login bug allowing people to log onto Unix hosts as root without entering a password (known publicly in September 1994) was solved in a fortnight under Linux, whereas IBM shipped AIX with the broken /bin/login for another couple of months (and didn’t tell customers although there was a fix available via FTP).

After-sales support for Linux, on the other hand, can be a good opportunity for small companies and startups. If you want to sell a Linux version of your product you might want to find such a company to outsource the after-sales support (see respective section). They might even help you with the porting.

Some companies selling big and expensive software packages offer completely installed systems, i.e. the hardware, the OS and their specific software. In the contract, you can restrict your warranty to specific kernel/library versions so that problems are unlikely to occur at a later time.

The GNU General Public License (GPL) and the GNU Library General Public License (GLPL) contain no restriction in terms of marketing products developed with GNU tools as the GNU C/C++ compiler. Anyway, the legal strength of the GPL was never checked at court – I doubt it can be enforced outside the U.S.A. If you violate the GPL, you may expect being sued by the FSF and being bashed publicly by the Linux community – the GPL has a strong ethic value. In case of doubts, you might want to contact the Free Software Foundation. A very comprehensive interpretation in German is printed in section 1.7 of Kai Petzke’s book ”Unix fuer jedermann – Einfuehrung in Linux”, Bernd-Michael Paschke Verlag, Berlin 1995, ISBN 3-929711-07-9.

In short terms: The GNU copyleft doesn’t affect your software development and marketing at all. Feel free to sell your products without source code and everything, just as with every other piece of commercial software. The GNU Public License says, if you put a piece of software under the GPL, or if you receive it as a piece of GPL’d software, you have to offer your customers the same rights concerning that software that you have (i.e. hand them the source code, or offer to deliver the source code for the cost of distribution for three years after delivery of the binary). The GPL does not force you to put all your software under the GPL, irrespective of whether the operating system it’s been written for is GPL’d itself. In this HOWTO, you’ll find an entire bunch of commercial software products none of which is GPL’d.

Pricing: There are three major philosophies. Some companies offer a full or limited version for a very low fee or even for free. Others consider the Linux version as valuable as those for other OS’s and offer it at a reasonable discount. Third, the price is equal to that of, say, the Solaris or the HP/UX version. Finding a reasonable price for your product may be somewhat more difficult than for other Unixes because there are many Linux users expecting to get everything for free. Others (seriously interested customers, i.e. business people and decision makers) know that nothing is for free. You’ll have to decide for the target customer.

A word concerning fast version shifts. This is clearly a prejudice. From the view of a software vendor, this is clearly not true. Linux is no moving target. It is true that Linus Torvalds <torvalds@cs.helsinki.fi> sometimes releases two or more kernel patches per week (or even more) in the hackers’ corner (odd minor kernel release numbers). However, if you found a kernel doing what it’s designed to do (version 2.0 being a good choice as of this writing), there’s no reason to upgrade. Leave everything as it is and limit your warranty to certain kernel releases. Selling complete systems helps here, too, because you can ensure the respective kernel and its device drivers work in perfect harmony with the computer, its peripherals and your application.

You might consider offering a statically linked Linux version too (as it is very popular for SCO due to some nasty bugs in several SCO runtime libraries) but please tell your customer about the additional memory requirement. However, it is important to mention that a dynamically linked application automatically satisfies all the requirements of any LGPL libraries that it uses, but a statically linked application does not. A statically linked application must be available in some relinkable form to satisfy the LGPL.

If you intend to offer a shrink-wrapped systems (i.e. a complete hardware/software installation) please avoid IDE harddisks. Those disks are a perfect performance bottleneck for a real-world multitasking OS if it comes to serious work. If some day EIDE offers busmaster DMA, this rule may change, though.

Two killer applications are needed under Linux: a good text processing system and a good spread sheet.

Copy protection is virtually impossible under Linux. Because the kernel and all the associated tools are available as source code, any experienced hacker can hook dongle or key card ioctl() calls. On a PC, you don’t have a machine/CPU id available. You don’t want to build a specialized kernel for a certain (and given) hardware configuration, not disclosing the corresponding source code; such a kernel would be a “derived work” according to the GPL, and doing so would clearly violate the GPL.

Last not least, don’t forget that university and college students who are the decision makers of tomorrow. If you offer them a reasonably priced product for their beloved operating system today (think of campus licenses), there’s little question which application they will choose tomorrow.

6 Linux for the Commercial Market?

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 iX Multiuser Multitasking Magazin Courtesy of iX Multiuser Multitasking Magazin!

People keep asking me “When is Linux going to be ready for the commercial market”. I guess the first thing to discuss is what is meant by ”commercial” in this context.

Some CD-ROM vendors have put the word ”commercial” in their name, only to have the technical people think their product is good only for use by banks and insurance companies. Other people look at their product with disdain and say that “Linux is not commercial quality”, because it is missing some feature they need, or they feel it is unstable (usually without ever trying it even one time).

To me, the word “commercial” has lost as much meaning in the marketplace as some of the other buzzwords:

• “Open” vs “Proprietary” (People are now saying UNIX is “Proprietary”)

• “Workstation” vs “PC” (What is a workstation, anyway?)

In the old days of computing the commercial market was banks, insurance companies and business-oriented facilities where the use of COBOL or RPG was the mainstream. The technical market was scientific, engineering and manufacturing where FORTRAN and assembly language was used. Somewhere along the way the term “commercial” seems to have gotten twisted around to mean “ready for the mass market”, versus “ready for hackers”.

For the purpose of this article I will take the second meaning, and address whether I think Linux is ready for commercial purposes rather than the hobbyist and hacker market, and ready for the mass market rather than limited markets.

For those of you who hate reading long articles, or who are short on time, let me give you my conclusion right now. Then you can go out and drink beer or other fun activities:

“Yes, Linux is ready for the commercial market...in some cases”.

In order for an operating system to be ready for the mass market it must have several attributes:

• have lots of applications

• be relatively easy to install

• have lots of applications

• be relatively easy to maintain • have lots of applications

• be relatively easy to use

• have lots of applications

• not crash (much)

• have lots of applications

• be economical

• have lots of applications

But you can eliminate all of these considerations in today’s mass market if only one thing is true:

You have lots of applications.

after all, there would not be 170,000,000 DOS systems in the world if any of the others had to be true.

I almost added that is has to be economical, but history has actually proven me wrong on that. If people added up the total cost of ownership, then Apple would certainly have won over the PC. But people ignore the human costs of someone else (or even themselves) beating their head against the wall trying to get something to work, or the system crashing repeatedly, or the fact that the one keystroke they can hit the easiest (through practice) is

             <CTRL><ALT><DEL>

In the old days people were content to spend several hundreds of dollars on a simple ASCII text editor, or deal with a simple spreadsheet. And it took an act of mangement to get them, with lots of Purchase Orders. Today, they want multi-media integrated with their operating system, and have all the applications available that their neighbor (or boss, or compatriot) has available on their system. And they want to get these applications easily, certainly no harder than to call up on the phone to order them through a catalog, or go down to their corner store to get them.

Now what causes this plethora of applications for an operating system? Ease of programming? Good software development tools? Features inside the operating system? Stability of the interfaces over time?

The answer is “none of these”. While all these attributes may help convince an application developer to port, the one overriding issue is volume of the operating system platform. Again, if MSDOS were compared to MacOS, or even to UNIX and volume were not taken into account, we know which two operating systems would have the most applications, and they would not be from Microsoft.

While it is true that several Linux vendors are working on getting these applications for the mass market (read this ” your mother and father”), the number of applications that run on Microsoft platforms have been estimated as high as 35,000. SunOS has an estimated 10,000 applications, with other ‘commercial UNIX” systems (including Solaris 2.x) much lower in number. It will take the Linux vendors a long time to get the number of applications necessary to hit the really large mass market, particularly if they did not depend on iBCS2 and DOS/Windows compatibility (which could supply a fair number of current applications), but depended on “native” Linux applications.

So applications are king (and queen) for the mass market, and installed base (volume) or the promise of explosive growth (volume) is the key to these. But is the mass market the only “commercial” market? The answer is “no”. The mass market is a subset (albeit very large one) of the commercial market. So let’s look at what the rest of the commercial market needs. We will look at this by segmenting the market into:

• turnkey systems

• large end-user customers

• specialized markets

6.1 Turnkey systems

When I speak of turnkey systems I typically mean a computer system that has one specific (or not so specific) application that runs on it. Examples of turnkey systems are point-of-sale terminals, reservation systems, CAD systems, etc. But in a larger sense, other applications such as Web servers, nameservers (such as BIND), etc. could also be considered “turnkey”, since they have only a few necessary programs that have to run on the system.

Usually turnkey systems are ones that an Independent Software Vendor (ISV) or Value Added Reseller (VAR) will chose a hardware system, an operating system, port an application to it, then duplicate that system 500 to 1000 times without change to the basic application. These ISVs and VARs will try to chose the lowest cost solution to fit their customer’s needs.

Linux is perfect for these types of applications. The operating system is stable enough for the developer to port their application and test the application fully. Once it is fully tested and stable, the entire package is “frozen” and duplicated any number of times for the end customer.

Since the operating system may be freely copied, and it runs on inexpensive hardware, their variable costs are minimal. Even a developer who is not familiar with the Linux system (so they need help getting it running on a platform) will quickly pay back the porting and system programmer costs they accrue by not paying dollar;200-dollar;500. per license for the operating system. Plus they have all the source code for the entire system, in case they run into trouble later on. You can buy a lot of Linux support for $200-$500K.

As I said before, I include Internet Service Providers (ISPs) as part of this “turnkey” environment, for both external internet and internal internet. Why overload your expensive, high-powered, highly complex general-purpose server to do Web serving when a smaller, simpler box can offload it? Why not run your NIS slaves on a Linux box? Or perhaps your BIND server?

In the early days of Digital UNIX (known then as DEC OSF/1) we did not have very many applications. In fact we had none. The marketing staff came to me with sad faces asking if it was possible to sell an operating system that had no applications. I invented a term called “Turbocharging”, which allowed a Digital UNIX system using the speed and power of the Alpha processor (as well as the throughput of our networking devices) to offload NIS, NFS, BIND and other services from people’s overloaded, slower SPARC machines. We also showed people how they could use the rsh(1) command to allow the Alpha to do a portion of their very CPU intensive processing while delivering the result back to the SPARCs on their desk. This allowed the SPARCs to work more on applications and less on the other “system administration” tasks that they were performing. We sold lots of Digital UNIX systems based solely on executing those tasks. Today, of course, Digital UNIX has a lot more applications, and particularly very large memory databases that are extremely fast. But the same principle applies. The database engine runs on the Alpha system, supplying data to the slower SPARC engines as a “Turbocharger”. I could see Linux systems headed in the same direction, following the same path.

6.2 Large end-user customers

Very large customers often have their own home-grown applications which they need to deploy across a wide network of people. Or they can have management dictate a certain suite of applications, which then can be ported to Linux. Since these customers are so large, their operating system costs are huge, and utilizing the savings using the Linux operating system they may completely cover the expenses of porting their software.

Or these very large customers may “influence” their layered product providers to port to the Linux platform. Finally, they may even change some of their computing habits (to use existing programs) if the cost savings are enough to warrant it.

Companies like Caldera are creating a suite of applications and approaching these very large customers to show them the operating system savings that they can achieve if they switch to Linux. While it is true that every application the customer could ever conceive of running may not run on Linux, by using the native applications, the iBCS2 applications, the DOSEMU applications, and applications that run under WABI, a nice suite of applications could be built to solve their needs.

6.3 Specialized markets

Finally there are what I call “specialized markets”. Markets that might buy Linux simply because it is Linux, and not because of the application suites that it provides.

In the education field there are three main markets:

• administrative

• “campus computing”

• computer science education

The administrative part is the “business” aspect of the market. They are looking for easy-to-use systems that can also handle complex administrative tasks that might cover a community the size of a small city.

The “campus computing” is the supply of computing power and service for majors of all types, web services and research into non-computer science (for example, molecular modeling) research.

Finally there is computer science education, both on the undergraduate and graduate level, as well as research into computer science.

While the administrative sub-market typically relies more on shrink-wrapped applications, the other two rely on them to a lesser extent (with the computer science education market relying the least). The other two markets can utilize a lot more of the freeware and shareware applications that are already ported to Linux. This gives them a very low-cost (from a software perspective) platform while allowing them to see and (often) modify the source code for the applications they use.

More importantly, in the computer science research area, the results of the research can be freely distributed to others working in the field, or even published as source code to illustrate the results. This can not be done with “commercial” operating systems.

Some universities are utilizing Linux more and more to run their campus. From a “commercial” standpoint, their needs are the same as many large businesses. Students graduating from college will know about Linux, and bring the word to their future employers.

Finally, there is the computer hobbyist and software developer market. I relate this market to the amateur radio market. In the amateur radio market the radio is often used to simply talk to other people, but at the same time the users investigate new ways of using radio, and improving it. Many electrical engineers started out as amateur radio users. So it can be with Linux, since for the first time both the prices of the hardware and the prices of the operating system source code are within the reach of mortal people.

In conclusion, I feel that Linux does have the items needed for several types of “commercial” uses:

• stability and quality

• low variable costs for turnkey applications

• explosive growth in volume to attract ISVs

What Linux really needs is for the “commercial” community to understand what is going on, and to embrace it where it will be useful. This will increase the volume numbers even more, which will attract more applications.

Along these lines I would like to “advertise” a joint effort of USENIX and Linux International to happen in January of 1997 in Anaheim, California of the United States. There will be a joint USENIX/Linux development conference, and while a certain part of the Linux conference will be oriented towards the development of the Linux operating system, the bulk of the conference will be oriented towards application developers and marketing people, to better understand the Linux operating system and how to sell their applications and services into the Linux market. We hope to show ISVs, VARs, resellers and distributors how they can make money by selling their applications and services on top of the Linux operating system.

To join in this effort, please view http://www.usenix.org/linux.html and encourage your VARs, ISVs and resellers (or even yourself!!) to attend. We call the conference ”USELINUX”, and we hope that we can prove to people at that event that it is time to just USE LINUX.

6.4 Biography

Jon “maddog” Hall is a Senior Leader in the Digital Equipment Corporation UNIX group. He has been in the computer industry for twenty-five years, UNIX for sixteen years and has guided the emergence of six operating systems, including Alpha Linux. He has an MS in Computer Science.

7. USELINUX Conference 16

7 USELINUX Conference

USELINUX: Linux Applications Development and Deployment Conference January 6-10, 1997 - Anaheim Marriott Hotel, Anaheim, California

Co-located with the USENIX Annual Technical Conference Co-sponsored by Linux International and the USENIX Association

========================================
Suggestions for Topics:  June 1, 1996
Submissions  Due :  Date:  July 1, 1996
Materials Due Date:  November 13, 1996
========================================

CONFERENCE CHAIR: Michael K. Johnson, Linux Journal

TECHNICAL TRACK COMMITTEE:
Michael K. Johnson, Chair, Linux Journal
Mark Bolzern, WorkGroup Solutions
Alan Cox, 3Com Remote Access Products
Jon ``maddog'' Hall, Esq., Digital Equipment Corporation
Lorrie LeJeune, O'Reilly and Associates
Dr. Tom Miller, North Carolina State University
Erik Troan, Red Hat Software
Dr. Greg Wettstein, Roger Maris Cancer Center
BUSINESS TRACK COMMITTEE:
Jon ``maddog'' Hall, Esq.,  Chair, Digital Equipment Corporation
Jonathan Eunice, lluminata, Inc.
Michael K. Johnson, Editor, Linux Journal
Lorrie LeJeune, O'Reilly and Associates
Bryan Sparks, President , Caldera, Inc.
Paul Winbauer, Director of Technical Programs, Avnet Computing
Bob Young, President, Red Hat Software, Inc.
==============================================
Tutorials:  January 6-7, 1997
Technical Sessions:  January 8-9, 1997
Business Sessions:  January 10, 1997
Keynote:  Wednesday, January 8, 1997
Birds-of-a-Feather Sessions: January 7-9, 1997
Vendor Exhibits:  January 8-9, 1997
Reception: January 8, 1997
==============================================

7.1 Overview

The Linux Applications Development and Deployment Conference (USELINUX) is aimed at three primary audiences: application developers porting or developing Linux applications, systems administrators charged with maintaining Linux systems, and business people who wish to bring Linux applications to market. Two technical tracks on January 8th and 9th will include separate components for developers and systems administrators. A business track devoted to explaining the dynamics of the Linux marketplace and how to partake of it will take place on January 10.

In addition to the three days of presentations and discussions, there will be two days of tutorials, Birds-of-a-Feather sessions, and Vendor Exhibits.

7.2 Tutorials

We are actively seeking proposals for half or full day tutorials on practical, technical aspects of using Linux. If you would like to present a tutorial, please contact the USENIX tutorial co-ordinator, Daniel V. Klein.

     Phone: 412.421.0285
     Fax: 412.421.2332
     Email: <dvk@usenix.org>

7.3 Technical track topics

The technical track will have three components:

• An application developers component focusing on topics helpful in porting and developing Linux applications. These include things like APIs, including both Linux’s levels of compliance with industry standard APIs and Linux-specific APIs, and capabilities that give extra functionality and/or convenience.

• A system administrators component to help sysadmins apply their skills to Linux administration, and also demonstrate some of the unique and useful features that can make their lives easier.

• A component for enthusiasts involved in developing the Linux operating system and environment.

7.4 Business track topics

The business track program will focus on obstacles and challenges in integrating Linux into a business. The target audience are software application developers, hardware vendors, service providers, large in-house application developers, and others who would like to know how to make their business more successful using Linux. This track will concentrate on business issues such as:

• The Linux Market: Who, What, Where, When and Why?

• Application Portfolios: What is available, what can be done?

• Marketing to the Linux Marketplace

• Channels: Retail, Resellers, Distributors, Integrators, OEMs, Service

• Licenses and Licensing: I don’t want to give away my application!!!

In addition to the meetings, there will be a compendium of information on CD-ROM for conference attendees. This will include copies of the slides, pointers to resources, white papers, lists of current resellers, user groups, etc.

7.5 How to submit a proposal

Suggestions for additional topics and areas where the topics can be expanded are welcome by June 1, 1996. The program committee will then prioritize the topics and create the final list of topics by July 1st. The program committee will solicit volunteers to work on the program, and will balance volunteers with program needs by that time.

An expanded electronic version of this announcement with greater detail, as well as the Call for Papers for the USENIX Annual Technical Conference, is available at the USENIX Web site at http://www.usenix.org.

Proposals for invited talks and panels should be received by July 1, 1996. We welcome submissions of a full paper or an extended abstract. Panel proposals should contain a list of names of potential panelists. Please send submissions to the conference program chairs via one of the following methods. All submissions will be acknowledged.

Please send comments, suggestions, and volunteering of time for a particular area of expertise (including a small bio of your experience) to the conference program chair via one of the following methods. Email is greatly preferred.

7.6 Preferred method

If you are submitting an idea for the technical track, send email to michael@usenix.org.

For the business program, send your ideas via email to maddog@usenix.org with the subject line:

     MADDOG, ANOTHER GREAT IDEA FOR THE *FABULOUS* USELINUX BUSINESS TRACK

7.7 Alternative method

Via postal address or fax to:

     Michael Johnson or Jon ``maddog'' Hall, Esq.
     USENIX Association
     2560 Ninth Street, Suite 215
     Berkeley  CA  94710
     Fax:  510-548-5738

7.8 Vendor exhibits

Vendors will demonstrate their products in a relaxed environment where attendees can discuss product features and services.

Vendors are invited to participate in the Vendor Exhibits. This is an excellent opportunity to receive feedback from our technically astute audience. If your company would like to display its products or services, please contact Cynthia Deno:

     Phone:  408-335-9445
     Email: <display@usenix.org>

7.9 Birds-Of-a-Feather sessions (BOFs)

BOFs are very informal, attendee-organized sessions held in the evenings by attendees interested in a particular topic. They may be scheduled on-site or in advance by contacting the USENIX Conference Office. Send email to conference@usenix.org or phone 714-588-8649.

7.10 About the USENIX Association

USENIX is the UNIX and Advanced Computing Systems Technical and Professional Association. Since 1975 the USENIX Association has brought together the community of engineers, system administrators, scientists, and technicians working on the cutting edge of the computing world.

The USENIX Conferences have become the essential meeting grounds for the presentation and discussion of the most advanced information on new developments in all aspects of advanced computing systems.

The USENIX Association and its members are dedicated to:

• problem-solving with a practical bias,

• fostering innovation and research that works,

• communicating rapidly the results of both research and innovation,

• providing a neutral forum for the exercise of critical thought and the airing of technical issues.

SAGE, a Special Technical Group within the USENIX Association, is dedicated to the recognition and advancement of system administration as a profession. To join SAGE, you must be a member of USENIX.

7.11 About Linux International

Linux International was established to promote the development and use of Linux. The people at Linux International know how good Linux is and want it to become an accepted competitor to products from even the largest computer companies. Among other things, it serves as a bridge between the dedicated and skilled community of developers, and the general world of commerce and industry.

Linux International is a non-profit organization run by volunteers, many of whom are high-profile Linux developers or activists.

For more information, please visit Linux International’s Web site at http://www.li.org or send mail to li@li.org.

7.12 Registration information

The complete program will be available in September 1996. To receive a registration package, please contact:

      USENIX Conference Office
      22672 Lambert Street, Suite 215
      Lake Forest, CA  92630
      Phone:  714.588.5738
      Fax:  714.588.9706
      URL:  http://www.usenix.org
      Email:  conference@usenix.org

Or you can send email to our mailserver at info@usenix.org. Your message should contain the line: “send linux conferences” and the program will be sent to you.

8 Linux vs. other free Unix flavors

Why port to Linux and not to one of the free BSD versions like 386BSD and its derivatives FreeBSD and NetBSD (short: *BSD)? Well, that depends on your application. If it was originally written for a SystemV-based system as AIX, HP/UX or Solaris 2, porting to Linux should be fairly easy. *BSD lacks certain SystemV features (such as the termio interface), and you’ll have to carefully check the amount of work necessary. On the other hand, FreeBSD and NetBSD are maintained by a closed developers’ team (as opposed to the open and sometimes chaotic and anarchic Linux developers’ community), and e.g. for FreeBSD there is only one so-called distribution available per FTP. You might want to mail info@FreeBSD.org or take a look at http://www.cdrom.com/titles/freebsd.html.

By the way, the FAQ for comp.unix.bsd.netbsd.announce concerning the *BSD flavors says (short excerpt):

        5.  Where BSD and POSIX differ, 386BSD conforms by default to
        BSD; Linux to POSIX.  Furthermore, while both run mostly GNU
        utilities, Linux tends toward the SysV flavor (e.g. init)
        where 386BSD sticks with the BSD style.  However, sources for
        different flavors of utilities are available for both, and
        both support compiler options which allow more BSD or more
        POSIX semantics.

And later it says:

        1) the 386BSD family started with BSD, and Linux started with
        POSIX.  NetBSD/FreeBSD/386BSD have been adding POSIX and System
        V compatibility, and Linux has been adding Berkeley and System
        V compatibility.  So there's a good deal of overlap.  But ...BSD
        is still a better choice if you want to program in a Berkeley
        environment and Linux if you want a POSIX environment.

9 Product Groups

This list is divided into several product groups to make finding a specific product more easy.

9.1 Databases

9.1.1 AccountFlex

Description:

AccountFlex is a powerfull full featured accounting system that includes modules for Order Entry, Inventory, Purchasing, A/R, A/P, G/L, Payroll, and Jobcost. Accountflex uses Infoflex, an SQL based 4GL lanquage that is compatible with Informix (see information in this HOWTO. hm).

Distribution Media:

Diskette, Tape.

Licensing Policy:

Source, runtime and unlimited runtimes are available. Demos and source are available for Linux, UNIX, DOS, or VMS.

OS Provisions:

Character Based Screen images.

Documentation:

Printed documentation.

Product Support:

Phone support charged by time usage. Onsite training and custom programming available.

Linux Support:

N/A Available Since:

October, 1994. Countries with Distribution:

USA. Price Range:

USD795 per module including source. Installed Base:

11 (as of date of entry. hm) Vendor:

          Infoflex Inc.
          840 Hinckley Road, Suite 107
          Burlingame, CA 94010, USA
          Phone: +1 (415) 697-6045
          Fax: +1 (415) 697-7696
          Contact: Gerard Menicucci

Last modified:

May 10, 1995.

9.1.2 CONZEPT 16

Description:

CONZEPT 16 is a complete software development system based on a relational database. It runs on different platforms and in heterogenous networks. A version with graphical user interface is available too. Client/Server technology optional. CONZEPT 16 is a high performance system, for which there are hardly limitations.

Distribution Media:

Diskettes. Licensing Policy:

Copy-protected. OS Provisions:

Kernel version 1.1.26; RAM: at least 2 MByte, recommended: 8 MByte, harddisk: about 3-4 MByte. Documentation:

German, english, training documentation and reference (700 pages each), Online help. Product Support:

Available (prices on request).

Linux Support:

N/A Extra Features and Add-Ons:

OEM-KIT, Recovery (prices on request). Available Since:

September 1, 1994. Countries with Distribution:

Worldwide. Price Range:

Upon request. Installed Base:

N/A Vendor:

          vectorsoft Gesellschaft fuer Datentechnik mbH
          Seligenstaedter Grund 2
          D-63150 Heusenstamm, Germany
          Phone:  +49 6104/6477
          Fax:    +49 6104/65250
          Mailbox: +49 6104/5022

Last modified:

April 6, 1995.

9.1.3 D-ISAM

Description:

Multikey B+ tree Isam File Handler. Follows the C-ISAM (Informix) file structure and function calls. Sold with Source.

Distribution Media:

5.25” and 3.5” Floppy disk. Licensing Policy:

Free distribution of the first 100 D-ISAM’d executables. Royalties apply after that. OS Provisions:

Independent. Documentation:

Documentation in english. Unix Style manual.

Product Support:

EMail, bbs, fax, telephone. First 90 days free. US$450/year for ongoing.

Linux Support:

N/A

Extra Features and Add-Ons:

Transaction Processing functions: US$295.00

Available Since:

June 1994

Countries with Distribution:

UK, Germany, USA, Canada (manufacturer).

Price Range:

Single Programmer License with Source: US$595, Transaction Processing Add-on: US$295, Additional Programmers: US$100, Runtime Licenses per 100: US$695 (after the first 100), Unlimited Runtime: US$6950. A free D-ISAM for Linux library is available for development, but not for commercial use, via FTP from ftp://ftp.wimsey.com/pub/linux/d-isam.

Installed Base:

Vendor:

     Byte Designs Ltd
     20568 - 32 Avenue
     Langley, BC V3A 4P5, Canada
     Phone:  +1 (604) 534-0722
     Fax:    +1 (604) 534-2601
     Contact: Heinz Wittenbecher
     E-mail: sales@byted.com

Last modified:

September 8, 1994.

9.1.4 Empress RDBMS

Description:

The port to Linux provides customers a choice of two levels of product: A) full function RDBMS B) single user Personal EMPRESS for Linux. A) EMPRESS RDBMS is a powerful application development tool which provides heterogeneous distributed processing capabilities. Its features include data manipulation options, integrity and recovery features, portability, and unlimited storage capacity. B) Personal EMPRESS for Linux is a low-cost, single-user version designed especially for the rapidly growing Linux users community. It includes Version 6.6 of the EMPRESS RDBMS and 4GL application generator; SQL, Dynamic SQL and host language C interface; and the EMPRESS GUI Builder for Motif. This product is available via catalog through the ACC Bookstore.

Distribution Media:

CD-ROM. Licensing Policy: OS Provisions:

Linux kernel versions 1.0.8, 1.1.72 and 1.2.1. Documentation:

A) printed manuals. B) CD-ROM. Product Support:

Technical support includes software updates, telephone support, internet access to Empress Technical notes, and a quarterly subscription to the users’ newsletter Empressions. A) Full technical support is provided with the full-function, multi-user version (EMPRESS RDBMS). B) Technical support is available at an additional cost of USD1000 annually (Personal EMPRESS for Linux).

Linux Support:

A) none. B) if purchased through the ACC Bookstore, can be bundled with Red Hat Linux support. Extra Features and Add-Ons:

EMPRESS RDBMS options include the EMPRESS ODBC Interface for allowing PC applications to access EMPRESS databases. Other features and functions include shared library and memory, and math library functions.

Available Since:

May 1, 1995. Countries with Distribution:

Empress products are available worldwide through our network of sales offices and distribution partners, as well as through the ACC Bookstore. For specific information please contact Empress International Operations: tel +1 (905) 513-8888, fax +1 (905) 513-1668.

Price Range:

A) Please contact vendor for pricing of EMPRESS RDBMS. B) Personal EMPRESS for Linux is available at a cost of USD80 in U.S. and Canada, and USD120 elsewhere. This includes tax and shipping.

Installed Base:

A) 10 copies. B) 250 copies. Vendor:

     A)
     Empress Software Inc.
     6401 Golden Triangle Drive
     Greenbelt, MD 20770, USA

     Phone:  +1 (301) 220-1919
     Fax:    +1 (301) 220-1997
     E-mail: sales@empress.com
     Contact: Jim Sweeney
     B)
     ACC Bookstore
     Phone:  +1 (800) 546-7274
     E-mail: bob@redhat.com
     Contact: Bob Young

Last modified:

June 14, 1996.

9.1.5 ESQLFlex

Description:

ESQLFlex is a low cost clone of the Informix-ESQL/C and Informix standard Engine products. ESQLFlex will allow developers to completely replace Informix-ESQL and Informix standard engine without having to modify their existing application. ESQLFlex enables developers to build, modify, and/or query databases using standard SQL calls from within ”C” programs.

Distribution Media:

Diskette, Tape Licensing Policy:

Source, runtime and unlimited runtimes are available. Demos and source are available for Linux, UNIX, DOS, or VMS.

OS Provisions:

Character Based Screen images. Documentation:

Printed documentation. Product Support:

Phone support charged by time usage. Onsite training and custom programming available. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

October, 1994.

Countries with Distribution:

USA, Spain, France, Germany. Price Range:

US $ 495. Installed Base:

35 (as of date of entry. hm). Vendor:

          Infoflex Inc.
          840 Hinckley Road, Suite 107
          Burlingame, CA 94010, USA
          Phone: +1 (415) 697-6045
          Fax: +1 (415) 697-7696
          Contact: Gerard Menicucci

Last modified:

May 10, 1995.

9.1.6 InfoFlex

Description:

InfoFlex is a complete 4GL that streamlines the design process with a consistent WYSIWYG approach to developing menus, screens and reports. Infoflex is compatable with the Informix database and is similar in syntax. Demos, source, and unlimited licenses are available for UNIX, DOS, or VMS.

Distribution Media:

Diskette, Tape. Licensing Policy:

Source, runtime and unlimited runtimes are available. Demos and source are available for Linux, UNIX, DOS, or VMS.

OS Provisions:

Character Based Screen images. Documentation:

Printed documentation. Product Support:

Phone support charged by time usage. Onsite training and custom programming available. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

October, 1994. Countries with Distribution:

USA, Spain, France, Germany. Price Range:

US $ 995. Installed Base:

18 (as of date of entry. hm). Vendor:

     Infoflex Inc.
     840 Hinckley Road, Suite 107
     Burlingame, CA 94010, USA
     Phone: +1 (415) 697-6045
     Fax: +1 (415) 697-7696
     Contact: Gerard Menicucci

Last modified:

May 10, 1995.

9.1.7 Just Logic/SQL Database Manager

Description:

The Just Logic/SQL Database Manager is a relational database system made to be used from C and C++ applications. It includes a complete set of libraries, utilities and a database engine. It comes with three programming interfaces: a C API interface, a C++ Class definitions and a C Precompiler. The client-server version permits Windows and Unix applications to access a remote server on Unix across TCP/IP.

Distribution Media:

Diskettes. Licensing Policy:

Unlimited runtime licence of applications. OS Provisions:

SLS 1.0, 400K RAM, plus 100K per user, and 1.2 Meg Hard Disk, plus 1 Meg for client-server.

Documentation:

User manual (175 pages). Product Support:

Phone, EMail, fax. Linux Support:

N/A Extra Features and Add-Ons:

Fully compatibles versions on SCO, BSDI, UnixWare, FreeBSD, DOS, Windows and OS/2. Available Since:

August, 1994. Countries with Distribution:

Sold to all countries. Distributors in US and Germany. Price Range:

The Linux release is priced at USD149 (USD219 with client-server). Installed Base:

250 (Linux release). Vendor:

          Just Logic Technologies Inc.
          P.O. Box 63050
          40 Commerce St.
          Nun's Island, Quebec H3E 1V6, Canada
          Phone: +1 (800) 267-6887 (toll free USA and Canada)
                 +1 (514) 761-6887
          Fax: +1 (514) 642-6480 (has voice-fax autodetection)
          E-mail: 71563.3370@CompuServe.COM
          WWW: http://www.pht.com/justlogic

Last modified:

September 08, 1994.

9.1.8 POET 2.1

Description:

POET 2.1 ODBMS for C++. Database functionality for C++ objects. Full support of encapsulation, inheritance, polymorphism and objectidentity. Two versions available : Personal Edition (Single/User 1 Developer) Professional Edition (Client/Server up to 4 Developer).

Distribution Media:

Diskettes. Licensing Policy:

SDK license from US$ 99. Runtime Licenses for commercial Products and for Public Domain or Share Products available.

OS Provisions:

Kernel version > 1.0; SDK needs 4 MBytes on harddisk. Documentation:

Handbook and tutorial (450+ pages). Product Support:

Hotline support, training please contact vendor. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

March, 1994. Countries with Distribution:

Many; please contact vendor. Price Range:

Personal Edition US$ 99 / DM 170. Professional Edition US$ 1990 / DM 3.990 (includes 1 year hotline support).

Installed Base:

234 copies as of July 1994 Vendor:

          POET Software GmbH
          Fossredder 12
          D-22359 Hamburg, Germany
          Phone:  +49 (0)40/609 90 18
          Fax:    +49 (0)40/603 98 51
          E-Mail: info@poet.de
          Contact: Detlef Meyer

Last modified:

September 19, 1994.

9.1.9 /rdb

Description:

/rdb is a RDBMS consisting of more than 125 shell level commands which read tables from the standard input and write tables to the standard output. Applications are typically written in shell scripts, mixing /rdb commands with ordinary system commands. A runtime library is also included.

Distribution Media:

3.5” diskettes. Licensing Policy:

/rdb is licensed by the CPU. Time stamped licenses are available to allow for hardware upgrades. An example of a full-text search using /rdb can be found at the Melody Shop at http://www.branch.com/ rsw.

OS Provisions:

Kernel version 0.99.12 or later; consumes about 3MB storage. Documentation:

230 page manual, both printed and on line. Some on-line help. Informational and error message tables provided for translation.

Product Support:

Free EMail support; maintenance costs 10% of license fee/year. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

October, 1993. Countries with Distribution:

US, Canada, South America, Australia, Germany, France, Ireland, Great Britain. Price Range:

USD149 qty 1 / CPU. Installed Base:

101+ Vendor:

     Revolutionary Software
     131 Rathburn Way
     Santa Cruz, CA 95062-1035
     Phone:  +1 (408) 429 6229
     Fax:    +1 (408) 427 0342
     E-mail: rdb@rsw.com
     Contact: Evan Schaffer

Last modified:

April 7, 1995.

9.1.10 Veritas

Description:

Veritas is a system for German drinks wholesalers (”Getränkegrosshändler”). It’s based on Onyx which is my 4gl and database project, which of course stays under the GNU Public License 2.0. The first version of Veritas was developed for PCOS Olivetti in 1992, later Xenix versions followed. The recent 3rd release features twelve years of experience with this kind of market, so you can expect to have nearly anything a drinks wholesaler needs, e.g. Pfand, Sektsteuer, Brauereiabrechung, Bruchvergütung, etc. Several servers can be linked via SLIP or UUCP and replicate transactions to share data. Clients are connected via ethernet to a local server and can run MS Windows or better Linux.

Distribution Media:

Directly installed on the customer’s hardware. Plus a Tape containing (I hope) all of the GNU’ish sources including Onyx and a bunch of different databases.

Licensing Policy:

Veritas is licensed according to a German translation of the GNU Public License 2.0, restricted in a way that any changes made must be mailed to the author, and that any customer must know that I’m the author and give support. Although the licence is very much like the GPL, I really recommend to purchase a copy if you prefer to get support.

OS Provisions:

Linux 1.0.9 (MCC), XFree86 2.1. Servers: 8-16 MB RAM, 200-500 MB harddisk, tape, modem. Clients: 8 MB RAM, 40 MB harddisk.

Documentation:

About 41 pages printed for Veritas in German, and about 43 pages online documentation for Onyx in English.

Product Support:

Retail price includes one week of training, and one year first class support over modem lines and D2 phone directly from the author. If you are interested in Onyx support, contact the author directly.

Linux Support:

Support for the installed sites is included in the installation price for one year. Extra Features and Add-Ons:

N/A Available Since:

The 3rd version of veritas is now in BETA in Bremerhaven. Countries with Distribution:

Only distributed in Germany because of the user interface language and a lot of German tax related features. If sold outside the northern part of Germany we need a VAR able to handle Linux, and sell and install the right hardware for a share.

Price Range:

Between 10,000 and 20,000 DM without hardware which depends on the number of servers and additional tayloring.

Installed Base: Vendor:

          Michael Koehne
          Loss Datensysteme
          Bremerstr. 117
          D-28816 Brinkum, Germany
          Phone: +49 (0)421/87 55 00
          V32bis: +49 (0)421/87 05 32
          Fax: +49 (0)421/87 55 51
          E-Mail: kraehe@nordwest.de (company)
                  kraehe@bakunin.north.de (private)

Last modified:

September 08, 1994.

9.1.11 Yard SQL

Description:

The YARD company offers 5 SQL products as follows:

• YARD-SQL – Relational SQL database server with compliance to X/Open XPG4 and ANSI SQL 92

• YARD-ESQLC – Embedded SQL for C.

• YARD-ODBC – ODBC interface for MS Windows clients • YARD-NET – Remote access to YARD databases via TCP/IP.

• YARD-X – Motif client for database access (No development tool) Distribution Media:

Floppy disk and streamer tapes Licensing Policy:

License number and activation key for each product and installation with user dependent licenses for the SQL-Server.

OS Provisions:

LINUX 1.0.9 and upwards compatible versions; Motif 1.2.3; RAM Usage: 1 MByte (minimum) for Shared Memory, 500 KByte per User; Disk Usage: YARD-SQL 7 MByte, YARD-ESQLC 1 MByte, YARD-ODBC 1,5 MByte, YARD-NET 1 MByte, YARD-X 5 MByte.

Documentation:

Printed 600 pages reference and users guide in German. English documentation will be available early 1995.

Product Support:

Release Update, Price 25% Version Update, Price 50% Technical support according to a special maintenance agreement (20% of total pricing). Training on request.

Linux Support:

None. Extra Features and Add-Ons:

All products also available for other widely distributed Unix systems (e.g. SCO UNIX, SPARC Solaris).

Available Since:

January, 1994. Countries with Distribution:

Current distribution in Germany through Distributor Ordix Software (info@ordix.de), Paderborn, Germany. Europeean/international distribution scheduled for early 1995.

Price Range:

YARD-SQL 590,00 (1 user) - 8.490,00 DEM (unlimited user) YARD-ESQLC 390,00 DEM YARD-ODBC 490,00 DEM YARD-NET 490,00 DEM YARD-X 990,00 DEM

Installed Base: Vendor:

     YARD Software GmbH
     Hansestr. 99
     D-51149 Koeln, Germany
     Phone:  +49 (0)22 03/45 71 30
     Fax:    +49 (0)22 03/45 71 31
     E-Mail: yard@yard.de
     Contact: Thomas Schonhoven (thomass@yard.de)

Last modified:

September 24, 1994.

9.2 Data Visualisation, CAD

CAD tools, renderers, OCR software, etc.

9.2.1 Ghostscript 3.x

Description:

Aladdin Ghostscript 3.n is a full PostScript Level 2 language interpreter. It can display PostScript files on the screen with X Windows, convert them to various raster formats (TIFF/F, GIF, PCX, PPM), and print them on many non-PostScript printers, such as the H-P inkjet and laser printers.

Distribution Media:

PC diskettes only, normally compressed with tar + gzip. A shareware version is available via FTP from tp://ftp.cs.wisc.edu/pub/ghost/aladdin for non-commercial use only.

Licensing Policy:

Redistribution in non-commercial contexts (i.e., if no money changes hands in connection with the distribution) is allowed without license. Any commercial distribution of any kind requires a commercial license, with a negotiated license fee. If you are a commercial organization and intend to distribute Aladdin Ghostscript in a manner for which you believe you don’t need a license, please check with us: ”in connection with” is only a shorthand for the terms of the actual non-commercial license.

OS Provisions:

Will run on any Linux version. Will display on any X11R4 or R5 compatible window system.

Documentation:

English, probably about 30 pages. On-line only (no printed doc). No on-line help.

Product Support:

No training available. Maintenance (bug fixing) and updates are available with commercial licenses and unavailable otherwise.

Linux Support:

N/A Extra Features and Add-Ons:

We offer commercial-quality sets of the 35 standard PostScript fonts at a price of $10 to end users and a much lower price to commercial licensees.

Available Since:

August 1994. Countries with Distribution:

We distribute worldwide. Our only current overseas distributor is in Germany. Price Range:

We will ship diskettes to end-users for a charge of $58 to cover service and materials costs. There is an additional charge for $50 for a custom configuration and of $20 (domestic) or $30 (international) for next-day mail. Commercial licensing prices are negotiated, and vary considerably.

Installed Base:

(L. Peter Deutsch estimates the number of users of the commercial version under Linux at about 12,000, as of Sep 23, 1994.)

Vendor:

     Aladdin Enterprises
     203 Santa Margarita Ave.
     Menlo Park, CA 94025, USA
     Phone:  +1 (415) 322-0103
     Fax:    +1 (415) 322-1734
     E-mail: ghost@aladdin.com

Last modified:

Aug 26, 1994

9.2.2 IDL

Description:

IDL is the pioneering software for data analysis, visualization, and application development. IDL’s features include flexible I/O, 2D plotting, 3D graphics, volume rendering, image processing, mathematics, statistics, a cross-platform GUI toolkit, plus a high-level, array-oriented programming language. Use IDL for visual data analysis, rapid prototyping, or application development. IDL programs are portable across Linux, Windows 3.1, Windows 95, Windows NT, Mac, PowerMac, UNIX and VMS.

Distribution Media:

CD-ROM and anonymous FTP <ftp://ftp.rsinc.com>.

Licensing Policy:

Free demo CD-ROM available or download via ftp. OS Provisions:

IDL supports linux kernel version 1.2.8 dynamically linked using Slackware 2.3. Statically linked version also included for use with Linux versions less than 1.2.8. 16 meg RAM minimum recommended. 50 meg disk space.

Documentation:

1600+ pages of documentation on paper and hypertext on-line help. Product Support:

Customer support service, maintenance/update service, training and consulting are available. In addition, a usenet newsgroup <news:comp.lang.idl-pvwave> is devoted to IDL.

Linux Support:

N/A Extra Features and Add-Ons:

HDF support, animation, volume slicer, high resolution mapping built-in. Available Since:

September 1995 Countries with Distribution:

(many; ask vendor). Price Range:

contact vendor/distributor Installed Base:

N/A Vendor:

     Research Systems, Inc.
     2995 Wilderness Place
     Boulder, Colorado, USA
     Phone:  +1 (303) 786-9900
     Fax:    +1 (303) 786-9909
     Email:  info@rsinc.com
     Distributor for Central Europe:
     CREASO GmbH
     Phone:  +49 81 05/250 55
     Email: 100137.2421@compuserve.com

Last modified:

September 14, 1995

9.2.3 MRJ Symbolic OCR

Description:

MRJ SOCR (Symbolic OCR) (TM) is an OCR program for Japanese text. SOCR reads TIFF files and recognizes Japanese text in scanned images. Output formats include PC, Mac, and Unix formats with Unicode, Shift-JIS, JIS, and EUC encodings. Also available for Sun OS.

Distribution Media:

3.5” Floppy Disks. Licensing Policy:

One SOCR license per workstation. OS Provisions:

Operating System: Linux 1.x, Window System: Motif run time library required, System RAM: 16 MB, Hard Disk: 6 MB.

Documentation:

Installation Guide and User Manual. All documentation is in English. Product Support:

Technical support available via telephone, EMail, and FAX. Linux Support:

The Linux version of SOCR is fully supported. See Product Support. Extra Features and Add-Ons:

Contact MRJ for information on OEM and software developer versions of SOCR. Available Since:

August 1994 Countries with Distribution:

Distribution from U.S. Available for export. Price Range:

Special Introductory Offer – $895 through 30 September 1994. Installed Base: Vendor:

     MRJ, Inc.
     10455 White Granite Drive
     Oakton, VA 22124, USA
     Phone:  +1 (703) 385-0700
     Fax:    +1 (703) 385-4637
     E-Mail: socr@mrj.com

Last modified:

Sep 08,1994

9.2.4 SISCAD-P 1.3-3

Description:

SISCAD-P is a 2D-CAD system that gives engineers a production implementation of new design technologies - parametrics, variational geometry, inference sketching, a fully customizable user interface, constraint-based modeling and feature-based modeling. Today there is only a German version of SISCAD-P available. Unfortunately we can not yet provide a SISCAD-P version in any foreign language but we are already working on an English version. The release date for this version is not certain yet.

Distribution Media:

available via FTP. On diskette and CD-ROM available from several Linux distributors.

Licensing Policy:

SISCAD-P is running on several Unix workstations and is available as shareware for Linux. The shareware version is a FULL featured package which is restricted to at maximum 10 parts with each 2000 objects. It is provided at no charge to the user for evaluation. With this concept, the user can test all features of SISCAD-P and has a chance to try before buying it. After a reasonable trial period, the user must make a registration payment and will receive a license code which will disable all limitations.

OS Provisions:

X Windows running under Linux. 8 MByte of memory minimum, 16 MByte or more recommended. Slackware Linux 0.19pl14 or newer. Fvwm or Mwm window manager recommended.

Documentation:

Online help and several documents in PostScript, DVI and ASCII format.

Product Support:

Product support for registered users via EMail, fax and BBS.

Linux Support:

N/A Available Since:

February, 1994. Countries with Distribution:

Germany (worldwide via FTP). Price Range:

DM 399 commercial users; DM 49 students version. Installed Base:

20 registered users as per Sep 22, 1994. Vendor:

          Staedtler Mars GmbH \& Co
          Geschaeftsbereich Informationssysteme - SIS
          Moosaeckerstrasse 3
          D-90427 Nuernberg, Germany
          Phone: +49 (0)911/3080-691
          Fax:   +49 (0)911/3080-692
          BBS:   +49 (0)911/3080-609, login: info, password: gast
          E-Mail: support@SIS.Staedtler.DE
          Contact: Helmrich Streitmatter

Last modified:

September 19, 1994.

9.2.5 TecPlot 6.0

Description:

Tecplot is a plotting program for visualizing and analyzing engineering and scientific data. The standard version includes XY, 2D and 3D-surface plotting. An optional extension (called 3DV) adds the capability to visualize 3D volumetric data. With Tecplot you don’t have to write a program, just input your data and start visualizing immediately. (for more information consult vendor. hm)

Distribution Media:

3.5-inch floppies, 150MB-QIC, 4mm DAT, 8mm tape, (CD-ROM available November 1994) Licensing Policy:

Single-computer license (node-locked) Evaluation copies are available. OS Provisions:

4 MByte is minimum RAM, required hard disk space is 20 MByte. Documentation:

Printed Users Manual 512 pages, partial online help info. English.

Product Support:

Update and technical support included with license for first three months and available for extra charge after that. Training classes are available; price is negotiable.

Linux Support:

N/A

Extra Features and Add-Ons:

3DV is an optional package of volumetric features available for an additional $300 USD.

Available Since:

August 1993

Countries with Distribution:

(many, please consult vendor. For Germany, see below. hm)

Price Range:

USD 995 to USD 1395 per Node-locked License

Installed Base:

Vendor:

          Amtec Engineering, Inc
          PO Box 3633
          Bellevue, WA 98009-3633, USA
          Phone:  +1 (206) 827-3304 (800-676-7568 in US/Canada)
          Fax:    +1 (206) 827-3989
          E-Mail: tecplot@amtec.com
          Contact: Tom Chan
          For Germany:
          GENIAS Software GmbH
          Erzgebirgstr. 2
          D-93073 Neutraubling, Germany
          Phone: +49 (0)94 01/92 00-11
          Fax: +49 (0)94 01/92 00-92
          Contact: Johannes Grawe
          E-mail: jo@genias.de

Last modified:

May 15, 1995.

9.2.6 XVScan Description:

Image scanning and manipulation software for HP ScanJet scanners. Distribution Media:

FTP or EMail preferred. 4mm DDS2 DAT tape or floppies (additional $US15 in US, $25 international).

Licensing Policy:

XVScan is based on XV and is not available in demo version due to licensing restrictions. Source is available.

OS Provisions:

Linux, tested with 1.2.x and 1.3.x, requires a recent generic SCSI driver support to be built in (no earlier than 1.1.79). Motif is not required. Any XFree version (X11R5, X11R6).

Documentation:

Online, WWW, English. Product Support:

Updates free for first year. Linux Support:

N/A Extra Features and Add-Ons:

Available for HP-UX and FreeBSD. Available Since:

May, 1995. Countries with Distribution:

Worldwide. Price Range:

$US50 for FTP or EMail shipping. Additional US $15 for media in the United States, US $25, internationally. Mastercard, Visa and Discover Credit Cards accepted.

Installed Base: Vendor:

          tummy.com, ltd.
          #807 300 S 16th Street
          Omaha, NE 68102
          Phone: (402) 344-4426

     Fax: (402) 341-7119
     Contact: Sean Reifschneider (jafo@tummy.com) or Evelyn
             Mitchell (efm@tummy.com)
     EMail: xvscan@tummy.com
     WWW: http://www.tummy.com/xvscan

Last modified:

June 14, 1996.

9.3 Development tools

Compilers, development environments and so on.

9.3.1 Amzi! Prolog & Logic Server

Description:

Full-featured, Edinburgh-standard (no data typing) system for embedding Prolog rule-based modules in C/C++ and other languages. Includes: compiler, listener, debugger, linker, EXE generator, and call-in/call-out Logic Server API.

Distribution Media:

3.5” diskettes. Licensing Policy:

The Professional Edition includes an unlimited, royalty-free license. The Personal Edition is limited to distributing applications for non-commercial use only.

OS Provisions:

About 250kb RAM for typical small application, 3-4mb disk. Documentation:

Printed 250pp manual. Various technotes on disk. Product Support:

Subscription Plus service provides automatic updates for a full year, $150. Custom development services available.

Linux Support:

No Linux system included. For Amzi! software, free tech support for registered users by phone, fax and e-mail.

Extra Features and Add-Ons: Available Since:

November 3, 1995.

Countries with Distribution:

Anywhere. Price Range:

Personal, $98. Professional, $298. Installed Base:

New Linux version. Existing Windows/DOS version has been in use since 1986. Vendor:

          Amzi! inc.
          40 Samuel Prescott Drive
          Stow, MA 01775
          USA
          Phone: +1 508 897 7332
          Fax: +1 508 897 2784
          Contact: Mary Kroening
          E-mail: info@amzi.com

Last modified:

June 13, 1996.

9.3.2 Basmark QuickBASIC

Description:

The Basmark QuickBASIC Compiler is a multi-user IBM-PC BASICA, MBASIC and Microsoft QuickBASIC Compiler designed to provide performance and consistency across a variety of machines (e.g. i386 and i486, Pentium, SPARC, RS/6000, HP PA-Risc) under Unix, AIX, SunOS, Linux, HP-UX, Xenix.

Distribution Media:

3.5” diskette Licensing Policy:

Per machine, no run-time restrictions. OS Provisions:

GNU GAS and LD must be installed. Documentation:

400+ page manual, release and installation notes. Product Support:

Updates available for $39.00 less shipping. Maintenance contracts available.

Linux Support:

N/A Extra Features and Add-Ons:

C-ISAM (Informix Inc.) Interface in C source code form. Cost $35.00. Available Since:

December 1993 Countries with Distribution:

to purchase from manufacturer directly. Price Range:

$195.00 (less shipping) Installed Base: Vendor:

          Basmark Corporation
          P.O. Box 40450
          Cleveland, OH 44140, USA
          Phone:  +1 (216) 871-8855
          Fax:    +1 (216) 871-9011
          E-Mail: jgo@ios.com  (for orders)
          Contact: Joseph O'Toole (for orders)

Last modified:

Sep 23, 1994

9.3.3 Clickables Description:

Clickables are CGI exectuables - easy to customize and easy to install. Available as single programs or total toolkit. Adding new modules weekly.

Distribution Media:

Download from Internet Website Licensing Policy: OS Provisions: Documentation:

Online: HTML page with demo, downloadable text version. Product Support:

Update service, online, phone.

Linux Support: Extra Features and Add-Ons:

Will develop custom applications or customize current modules. Available Since:

December 15, 1995. Countries with Distribution:

Worldwide through the Internet. Price Range:

Single modules as low as US $15, total toolkit is $95. Installed Base: Vendor:

          Clickable Software
          26 Medway, Suite 13
          San Rafael, CA 94901
          USA
          Phone: 415.456.5582
          Fax: 415.456.4018 Fax
          Contact:
          EMail: rayc@catch22.com
          WWW: http://www.catch22.com/clickables

Last modified:

June 14, 1996.

9.3.4 CODINE Description:

CODINE Job Management System – A Job-Queueing System that allows optimal utilization of a heterogeneous workstation cluster. The system features static and dynamic loadbalancing, checkpointing, support for parallel programs etc. It is available for SUN, HP, IBM, SGI, CRAY, CONVEC, DEC and LINUX

Distribution Media:

QIC tape, DAT tape, ftp Licensing Policy:

free demo available at ftp://129.69.18.15/pub/info/producer/genias/codine OS Provisions:

The GUI requires Motif, however the GUI is not necessary to use the system. A command line interface is available. RAM: 1,5 MB for the master daemon; HD: some Megs.

Documentation:

Manual in English available (also via ftp in PS format) Product Support:

GENIAS offers training and support for the CODINE system. Please ask for prices. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

1994 Countries with Distribution:

Switzerland, Sweden, Taiwan, USA, Italy Price Range:

N/A Installed Base:

some 20. LINUX PC’s are mostly only add-ons to a larger cluster. Vendor:

          GENIAS Software GmbH
          Erzgebirgstr. 2
          D-93073 Neutraubling, Germany
          Tel.: 09401/9200-11, Fax: 09401/9200-92
          Phone: +49 94 01/92 00-11
          Fax: +49 94 01/92 00-92
          Contact: Johannes Grawe
          E-mail: jo@genias.de

Last modified:

May 10, 1995

9.3.5 Distributed Interface Object Server System (DIOSS)

Description:

DIOSS is a high-level C Language API to an RPC based daemon that provides Motif 2.0 interfaces. It includes an Interface Builder Tool (DiossLive) and a Distributed Database Daemon. All daemons are freely distributed.

Distribution Media:

CD/ROM and Internet Download (with license password).

Licensing Policy:

Free demo at via FTP from ftp://in-tech.in-tech.com. Linux Personal Edition - can only be used for freeware or shareware. Linux Corporate Edition - can be used for full commercial development.

OS Provisions:

Linux 1.2.7 and greater. Requires Portmapper. SunOS and Solaris versions available. Documentation:

English language documentation English language on-line (Home Page) code and configuration examples.

Product Support:

1 year free maintenance upgrades (with license password). Free Home Page FAQ support free. Email support. All DIOSS servers are freely distributable.

Linux Support:

we will offer a Home Page Linux (and SunOS and Solaris) system configuration and setup FAQ - in development) wee supply a link to the Linux Portmapper source

Extra Features and Add-Ons:

DiossWebDB - a set of CGI programs that allow HTML browsers to make SQL queries against our database daemon: USD49 Linux, USD500 SunOS and Solaris. The Database daemon is free.

Available Since:

DIOSS has been available since August 1, 1995. DiossWebDB is available Sept. 1, 1995.

Countries with Distribution:

We target USA, Europe, and Asia. Price Range:

DIOSS: Linux Personal Edition (USD249 - Introductory: USD99 till November). Linux Corporate Edition (USD995). SunOS and Solaris (USD2495). DiossWebDB: Linux (USD49). SunOS or Solaris (USD495).

Installed Base:

N/A Vendor:

     DIOSS Corp.
     5205 Leesburg Pike 1200
     Falls Church, VA 22041, USA
     Phone:  +1 (703) 671-0706 (U.S. and Canada: 1 (800) 337-3839)
     Fax:    +1 (703) 671-5734
     Email:  toneil@in-tech.com
     Contact person: Tom ONeil

Last modified:

September 14, 1995

9.3.6 Discussion

Description:

C++ CGI with back end database supporting threaded discussions. Distribution Media:

FTP and WWW download. Licensing Policy:

Free for educational and charitable nonprofit use. Demo version available off of our web pages at http://www.pennant.com/aex.

OS Provisions:

Available currently in a.out format, has run on kernels from 1.1.58 onwards. Documentation:

Online help available. Product Support:

Currently in beta. Web based technical support discussion available. Buyers of the various betas updated at no charge to final release version (1.x).

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

Novemeber 14, 1995. Countries with Distribution:

Demo has been downloaded all over the world. Price Range:

$50.00 for one discussion system. $20.00 per additional discussion (Linux pricing only). Installed Base:

N/A Vendor:

          Pennant Media
          1411 Lincolnshire Dr.
          Champaign, IL 61821
          Phone: (630) 867-7291
          Fax:
          Contact: Nick Nafpliotis
          EMail: aex@pennant.com
          WWW: http://www.pennant.com/aex

Last modified:

June 14, 1996.

9.3.7 EZ Clone Description:

Software for programming handheld and mobile amateur radios. Distribution Media:

Diskette. Licensing Policy: OS Provisions:

Plain Linux. Documentation:

Printed reference manual. Product Support:

Free updates for a full year. Linux Support: Extra Features and Add-Ons: Available Since:

January 1996. Countries with Distribution:

Worldwide shipping from the US. Price Range:

$40.00 USD. Installed Base: Vendor:

     acsys
     120 Longview Drive
     Los Alamos NM  87544
     USA
     Phone: +1 505 672 4003
     Fax: +1 505 672 4103
     Contact:
     EMail: ezclone@acsys.com
     WWW: http://www.acsys.com

Last modified:

June 14, 1996.

9.3.8 Dynace

Description:

Dynace (pronounced like “dynasty” without the “t”) is a preprocessor, include files and a library which extends the C language with advanced object oriented capabilities, automatic garbage collection and multiple threads. Dynace is designed to solve many of the problems associated with C++ while being easier to learn and containing more flexable object oriented facilities. Dynace is able to add facilities previously only available in languages such as Smalltalk and CLOS without all the overhead normally associated with those environments.

Distribution Media:

3.5” diskettes Licensing Policy:

One license per programmer; Applications are royalty-free Info and DEMO on http://www.everest.com/algorithms or ftp://ftp.everest.com/pub/dynace.

OS Provisions:

N/A Documentation:

310 pages printed Product Support:

N/A Linux Support:

N/A Extra Features and Add-Ons:

Complete Windows development system Available Since:

December 1993 Countries with Distribution:

N/A Price Range:

Developer Edition (manual, libs & executables, some source): USD249; Source Edition (includes full source): USD599

Installed Base: Vendor:

          Algorithms Corporation
          3020 Liberty Hills Drive
          Franklin, TN 37064, USA
          Phone: +1 (800) 566-8991 or +1 (615) 791-1636
          Fax:   +1 (516)791-7736
          Contact: Blake McBride
          E-mail: blake@edge.net

Last modified:

June 9, 1995

9.3.9 Eiffel-S

Description:

Eiffel Compiler for Linux. Distribution Media:

CD-ROM, diskettes or FTP. Licensing Policy:

Version 1.3 Shareware, available from the Sunsite FTP archive. Version 2.0 commercial product. OS Provisions:

At least 8 MByte RAM and GCC 2.6.3 necessary. Documentation:

Included into delivery as ASCII text. Printed documentation available for extra fee. Product Support:

Available for the commercial version (between 300 DM and 1500 DM; dependent on license type) Training for Eiffel - 3 Days, 1680 DM.

Linux Support:

Extra Features and Add-Ons: Available Since:

1993. Countries with Distribution: Price Range:

Shareware registration is 80 DM. Actual version 2.0: personal license 345 DM, commercial license 790 DM.

Installed Base:

100. Vendor:

          SIG Computer GmbH
          D 35619 Braunfels
          Germany
          Phone: +49 6472 2096
          Fax: +49 6472 7213
          Contact:
          EMail: info@eiffel.de
          WWW: http://www.wwww.com/sig

Last modified:

June 14, 1996.

9.3.10 Finesse Description:

OSF/Motif GUI for shell scripts. Distribution Media:

Free full featured developer version available via FTP from

ftp://ftp.science-computing.uni-tuebingen.de/pub/finesse-light or floppy. Licensing Policy:

Finesse is a fully commercial product for other Unixes where nodelocked licenses are possible. The Linux version is freely available via FTP from ftp://ftp.science-computing.uni-tuebingen.de/pub/finesse-light/. On Floppy medium: DM 90.

OS Provisions:

Kernel 1.2, Motif-1.2.4, XFree86-2.1.1 libs, 16 MB RAM min Documentation:

Postscript File.

Product Support:

No support for free version. Linux Support:

No support for free version. Extra Features and Add-Ons:

Not appropriate Available Since:

Linux Version since May 1995. Countries with Distribution:

World. Price Range:

Free Demo. Commercial version with support on request. Installed Base:

N/A Vendor:

          science + computing GmbH
          Hagellocher Weg 71
          D-72070 Tuebingen, Germany
          Phone:  +49 (0) 7071/9457-0
          Fax:    +49 (0) 7071/9457-27
          E-mail: info@science-computing.uni-tuebingen.de
          WWW:    http://www.science-computing.uni-tuebingen.de/
          Contact: Olaf Flebbe

Last modified:

August 15,1995

9.3.11 Flagship

Description:

CA-Clipper5, Fox, dBase & beyond for Unix. XBase 4GL applications development system & database. Superset of CA-Clipper. Can be used to port.

Distribution Media:

3 3.5” Diskettes, other media extra charge. Licensing Policy:

No royalty or licensing for software you create and distribute. For more information available via FTP from ftp://ftp.wgs.com/pub2/wgs/Filelist.

OS Provisions:

Linux 0.99 and later, has been tested on many. Requires C compiler, libraries and linker. Make and other utilities useful. 4 MByte RAM & 5 MByte Hard disk needed to run. More is better.

Documentation:

1200 page printed manual extra charge, same manual online is shipped as part of the product.

Product Support:

Support via phone, email, fax (email preferred), unlimited for 30 days, chargable thereafter. Maintenance agreements including discounts on upgrades are available. All of this is country dependent.

Linux Support:

Country Dependent

Extra Features and Add-Ons:

FS2 Toolbox, FoxKit, and others. Country dependent.

Available Since:

July 1994

Countries with Distribution:

We sell to all countries, and have regional distributors in most major areas, and on all continents.

Price Range:

Country dependent. North America: Single user US$199, Unlimited User US$499 until Nov 15, 1994. US$499 Single & US$999 Unlimited after that.

Installed Base:

Vendor:

     multisoft Datentechnik GmbH
     PO Box 312
     D-82027 Gruenwald, Germany
     Phone:  +49 (0)89/6417904
     Fax:    +49 (0)89/6412974
     E-Mail: 100031.267@compuserve.com
     Contact:  Dorte Balek
     North America:
     WorkGroup Solutions, Inc
     PO Box 460190
     Aurora, CO 80046-0190, USA
     Phone:  +1 (303) 699-7470
     Fax:    +1 (303) 699-2793
     E-Mail: info@wgs.com
     Contact: Virginia Lane

Last modified:

Sep 19, 1994

9.3.12 Gen/X Description:

An X GUI/Application development system with real-time data capabilities. Little or no X experience required.

Distribution Media:

Internet download, diskettes, tapes. Licensing Policy:

30 day evaluation available. One license per host or site options. OS Provisions:

Linux 1.2.13, No Motif, straight Xlib, Recom. 8M RAM. 8MB HD. At least 800x640 screen resolution. Documentation:

Printed documentation (600 pages), English, downloadable MS Word docs (1 disk). Product Support:

Update service, maintenance, training ($200/hd) and unlimited help. Linux Support: Extra Features and Add-Ons:

Real time package (US $99), Customizable widgets and dyn objects free. Available Since:

January 12, 1996. Countries with Distribution:

USA. Price Range:

$490 Development, $99 Real-time. Installed Base: Vendor:

          AgeTek, Inc
          2308 North Mission Circle. Friendswood, TX. 77546
          USA
          Phone: 713-996-6096
          Fax: 713-996-6096

     Contact: Anthony Correll
     EMail: acorrell@phoenix.net
     WWW: http://www.agetek.com

Last modified:

June 14, 1996.

9.3.13 ICC11

Description:

ICC11 is a full featured C Compiler for the HC11 microcontrollers: including floating point support, interspersed C and assembly listing, and much more. It’s available in native Linux version, as well as DOS and OS2 versions.

Distribution Media:

It’s available on 3.5” diskette. Latest upgrades are available via FTP, usually for free or for a minimal charge.

Licensing Policy:

There is an early version of the compiler available for free use in ftp://ftp.netcom.com/pub/im/imagecft/icc11/free

OS Provisions:

There is no special hardware or software requirement to run ICC11. Documentation:

The product comes with a 100 page manual. Product Support:

Preferred technical support via e-mail. Send mail to imagecft@netcom.com. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

Feb 95. Version 2 with floating point support available now. Countries with Distribution:

N/A Price Range:

USD45 plus USD5 S&H for U.S.A. and Canada, USD10 S&H for elsewhere.

Installed Base:

We have over 400 customers worldwide. Vendor:

          ImageCraft
          P.O. Box 64226
          Sunnyvale, CA 94088-4226, USA
          Phone/Fax: +1 (408) 749-0702
          Contact: Richard Man or Christina Willrich
          E-mail: imagecft@netcom.com

Last modified:

April 10, 1995

9.3.14 INSURE++

Description:

Insure++ (Formerly Insight++) is the most thorough automatic runtime debugger available. Insure++ identifies compile-time, runtime, and algorithmic errors, including memory overwrites, pointer errors, and errors in system calls, C/ C++ libraries, third-party libraries and I/O statements. Insure++ also includes data visualization tools and coverage analysis tools.

Distribution Media:

DAT tape, QIC tape Licensing Policy:

free demo at ftp://129.69.18.15/pub/info/producer/genias/insure. OS Provisions:

N/A Documentation:

N/A Product Support:

N/A Linux Support:

N/A Extra Features and Add-Ons: Available Since:

Countries with Distribution:

USA, Sweden, Netherlands

Price Range:

DEM 895 (with 25% educational discount)

Installed Base:

few

Vendor:

     ParaSoft Corporation
     2031 S. Myrtle Ave
     Monrovia, CA  91016
     Phone: 818-305-0041
     Fax: 818-305-9048
     E-mail: info@parasoft.com
     GENIAS Software GmbH
     Erzgebirgstr. 2
     D-93073 Neutraubling, Germany
     Tel.: 09401/9200-11, Fax: 09401/9200-92
     Phone: +49 94 01/92 00-11
     Fax: +49 94 01/92 00-92
     Contact: Johannes Grawe
     E-mail: jo@genias.de

Last modified:

May 10, 1995

9.3.15 INTERACTER

Description:

INTERACTER is a portable user- interface and graphics subroutine library for Fortran software developers. INTERACTER-based programs are portable to a wide range of Fortran compilers running on Unix, VMS, DOS and Windows. The Linux version supports f2c and g77. Functions include menus, text windows, forms, graphics primitives, presentation graphics, hardcopy (many formats) and operating system interface.

Distribution Media:

3.5” tar diskettes or 4mm DAT.

Licensing Policy:

INTERACTER-based programs can be distributed royalty free. Linux demo program available at http://www.demon.co.uk/issltd OS Provisions:

Linux version supports X Windows (via Xlib), graphics terminals or Linux console (colour text-mode). Requires g77 or f2c.

Documentation:

Two-volume English language printed manual.

Product Support:

Free lifetime technical support. New releases every six months. Chargeable software updates.

Linux Support:

Same as for all other platforms (see above).

Extra Features and Add-Ons:

Optional INTERACTER Toolkit provides form designer, graphics file viewer, on-line help, hardcopy post-processor and set-up file editor.

Available Since:

October, 1995.

Countries with Distribution:

Distributors in North America, Europe and Australia.

Price Range:

Linux version: US $675 (UK pounds 450).

Installed Base:

Vendor:

          Interactive Software Services Ltd.
          Westwood House
          Littleton Drive
          Huntington
          Staffs WS12 4TS
          United Kingdom
          Phone: +44 (0) 1543 503611
          Fax: +44 (0) 1543 574566
          Contact:
          EMail: support@issltd.demon.co.uk
          WWW: http://www.demon.co.uk/issltd

Last modified:

June 14, 1996.

9.3.16 ISE Eiffel 3

Description:

ISE Eiffel 3 provides a powerful and user-friendly O-O programming environment designed for large, complex systems. It is an integrated GUI workbench consisting of a variety of Eiffel-based components: EiffelBench melting-ice workbench, EiffelBuild interface builder and application generator, EiffelVision graphics and GUI library, and EiffelBase basic libraries.

Distribution Media:

1/4 inch Tape, DAT.

Licensing Policy:

Each hardware platform type requires a different runtime. The standard delivery of the EiffelBench comes with the runtime for the given platform type; additional runtimes may be purchased for cross-development. ISE charges run-time license fees, only for copyrighted commercial products built using ISE’s technology. For further information contact vendor.

OS Provisions:

Motif required

Documentation:

Documentation including the Prentice Hall books "Eiffel: The Language, Reusable Software: The Base Object-Oriented Component Libraries" and "An Object-Oriented Environment: Principles & Practices".

Product Support:

The purchase prices include free technical support for three months. The price of technical support per year is 30% (payable in advance) of the standard software purchase and provides assistance with problems encountered in the use of the software. Support is provided by electronic mail, fax or telephone.

Linux Support:

N/A

Extra Features and Add-Ons:

A special license for students and individuals of ISE Eiffel on Linux includes the full graphical environment.

Available Since:

June 1994

Countries with Distribution:

(many individual distributors worldwide; too many to name them here. please consult vendor. hm)

Price Range:

US$295.00 plus shipping & handling. USA: $15.00 for UPS Ground, $20.00 for UPS 2nd Day Air; Canada: $35.00 for UPS 2nd Day Air; Other: $60.00 for UPS International Air

Installed Base: Vendor:

     Interactive Software Engineering, Inc.
     270 Storke Road, Suite 7
     Goleta, CA 93117, USA
     Phone:  +1 (805) 685-1006
     Fax:    +1 (805) 685-6869
     E-mail: queries@eiffel.com

Last modified:

Sep 09, 1994

9.3.17 Metacard

Description:

MetaCard is a hypermedia/Rapid Application Development environment for X11/Unix workstations that is compatible with Apple Corp.’s HyperCard. MetaCard can be used by programmers and sophisticated end users to build Motif applications and hypermedia documents using a powerful, direct manipulation editor and a simple scripting language. Stacks developed with MetaCard are portable among all supported platforms (14 for release 1.3) and can be distributed with the MetaCard engine without licensing fees or royalties.

Distribution Media:

Anonymous FTP and 3.5” diskettes. Licensing Policy:

MetaCard can be licensed to a single, named individual, but can be used on any machine or combination of machines by that user. Multiple user packages are also available. The save-disabled distribution is available via anonymous FTP from ftp://ftp.metacard.com/MetaCard and ftp://ftp.uu.net/vendor/MetaCard.

OS Provisions:

The 1.4 release was built on Linux 0.99pl14. 8 MByte RAM minimum, 16MByte recommended. 800x600 minimum display resolution. Requires about 5 MByte disk space.

Documentation:

Complete on-line documentation is supplied. Printed documentation is an extra-cost option. Product Support:

Free email technical support. Free upgrades during the 1.X series of releases.

Linux Support:

N/A

Extra Features and Add-Ons:

A library version of MetaCard that can be linked directly to C programs (Embedded MetaCard) is available at extra cost.

Available Since:

February 1994

Countries with Distribution:

Direct sales world wide.

Price Range:

$247.50. $495 for a license that works on Linux and all Unix platforms.

Installed Base:

Vendor:

     MetaCard Corporation
     4710 Shoup pl.
     Boulder, CO 80303, USA
     Phone:  +1 (303) 447-3936
     Fax:    +1 (303) 499-9855
     E-Mail: info@metacard.com
     Contact: Scott Raney

Last modified:

Sep 20, 1994

9.3.18 Mjolner BETA System

Description:

The Mjolner BETA System is a software development environment supporting object-oriented programming in the BETA programming language. BETA is uniquely expressive and orthogonal amongst object-oriented languages. BETA unifies just about every abstraction mechanism, including class, procedure, function, coroutine, process, and exception, into the ultimate abstraction mechanism: the pattern. In addition it has general block structure, strong typing, whole/part-objects, and concurrency. Compiler: Native code generation, garbage collection, separate compilation, C and assembler interface. System: Persistence, distribution, comprehensive libraries, platform- independent GUI application framework (Unix, Mac, Windows, NT), metaprogramming system, graphical system, hyperstructure editor. Among the unique characteristics of the Mjolner BETA System are that the GUI-applications run on Motif, Windows NT and Macintosh without any change and with the native look and feel; it is easy to write distributed applications and it is easy to save and restore objects from a persistent store.

The Mjolner BETA System is being used successfully for teaching both first grade and second grade students at several universities.

Distribution Media:

3.5” Floppy Diskettes, FTP.

Licensing Policy:

The normal version requires a licence per CPU. Server licences are available at reduced cost. Educational licences do not permit the use of the system for commercial purposes, but permit free distribution of Personal Editions (see below) for students. A Personal Edition containing compiler, the basic libraries, persistence, X frameworks is available. This version may not be used for commercial purposes and/or teaching.

OS Provisions:

Linux 1.0 or later, RAM: 4Mb (8 to 16 recommended), XFree86, OSF/Motif 1.1 or later (for MotifEnv programs only; not required to run compiler), Programs needed: (g)as, ld, make.

Documentation:

Full system includes 1000 pages of printed documentation. FTP customers have access to all documentation in PostScript format. The Personal Edition includes 200 pages of printed documentation, and remaining documentation in PostScript.

Product Support:

Upgrade: 30% of the prices above. Hot-line support: USD500 per month per CPU license. E-mail support available. Ask for details.

Linux Support:

N/A

Extra Features and Add-Ons:

     Persistent store                                        USD  95
     X frameworks, including Xtenv, AwEnv and MotifEnv       USD 195
     Bifrost, the interactive graphics system                USD 195
     Yggdrasil, the metaprogramming system                   USD 145
     Sif, the hyper structure editor                         USD 195

Available Since:

August 1994

Countries with Distribution:

The Mjolner BETA System for Linux must be purchased from manufacturer directly.

Price Range:

Basic system is USD195. Add-ons ranges from USD95 to USD195. Full system available for USD820. A Personal Edition (non-profit) containing compiler, the basic libraries, persistence, X frameworks is available for USD60.

Installed Base:

N/A

Vendor:

     Mjolner Informatics Aps
     Science Park Aarhus
     Gustav Wieds Vej 10
     DK-8000 Aarhus C, Denmark
     Phone:  +45 86 20 20 00
     Fax:    +4) 86 20 12 22
     Contact: Peter Andersen (Peter.Andersen@mjolner.dk)
     E-mail: info@mjolner.dk, usergroup-request@mjolner.dk,
             sales@mjolner.dk

Last modified:

August 15, 1995

9.3.19 SEDIT, S/REXX

Description:

SEDIT is a powerful UNIX (tm) text editor patterned after IBM’s XEDIT editor. It operates with a GUI under X windows or in character mode from a tty device. S/REXX is a full UNIX implementation of IBM’s SAA procedural language except that the numeric digit specification is limited to 15 digits. S/REXX functions as an imbedded macro langauge for SEDIT as well as providing a powerful modern programming language alternative to shell scripting languages. SEDIT and S/REXX may be purchased in a bundle or individually. See http://www.portal.com/ sedit or http://www.sedit.com/sedit for more information including introductory pricing for Linux, etc. The ftp directory includes a text file flattened from the above web pages for users wouth WWW capability or mail us at sedit@shell.portal.com.

Distribution Media:

The Linux version is available for ftp download or on diskettes.

Licensing Policy:

SEDIT and S/REXX are licensed products. See the above URLs for details. For demo purposes, a short term license key will be provided on request to anyone who obtains the product via ftp.

OS Provisions:

We believe these products will work with any recent stable Linux kernel but testing has been limited to two: circa 1.0.5 and 1.1.49. Motif is fully bound with SEDIT and SEDIT is distributed with XFree86 3.1 shared libraries for use if your system is not at that level.

Documentation:

Documentation is furnished in PostScript form, in extensive on-line help files, and in optional printed manuals.

Product Support:

Technical support is available for all licensed users. For maintenance, see the above URLs. Linux Support:

n/a Extra Features and Add-Ons:

A REXX interactive graphical debugger is available as an optional feature as are printed manuals, physical media, and on going maintenance. See the above URLs.

Available Since:

The beta Linux version has been available since about May 1, 1995. The poduct version is anticiapted about July 1.

Countries with Distribution:

SEDIT and S/REXX are available worldwide. The URLs should be consulted for current information. Price Range:

Ranges from USD99 (introductory 2 user license without media or printed documentation) to USD440 (regular price for the 4 user version of SEDIT with S/REXX and all hardcopy manuals).

Installed Base:

N/A Vendor:

     Benaroya
     31 Rue de Constantinople
     F-75008 Paris, France
     Phone: +33 1 47 22 22 13
     Fax:   +33 1 47 22 06 17
     Contact:  Robert Benaroya
     E-mail: sedit@shell.portal.com

Last modified:

June 9, 1995

9.3.20 Smalltalk/X language system

Description:

Smalltalk language implementation with integrated development environment and smalltalk-to-C compiler. Includes incremental compiler, symbolic debugger, source code browser, class library (+600 classes), batch compilers for high performance machine code, inline C capability, lightweight processes, 3D graphics and much more.

Distribution Media:

Diskettes, CD-ROM, tapes or others. Licensing Policy:

Free demo version for non-commercial use. Full versions available. A free demo or shareware version is available at ftp://ftp.informatik.uni-stuttgart.de/pub/stx.

OS Provisions:

Linux 0.99 - 1.2; 16MB RAM minimum. Documentation:

Online; HTML. Documentation also available online via the WWW at

http://www.informatik.uni-stuttgart.de/stx/stx.html. Product Support:

Update service via FTP. Maintenance, training, consulting upon request. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

1994. Countries with Distribution:

USA, Germany. Price Range:

Students, Educational institutes (non commercial): US$ 150. Others (non commercial): US$ 300. Full commercial: US$ 3000. No per-runtime royalties.

Installed Base: Vendor:

          DevCon, Development \& Consulting
          Kirchheimerstr. 87
          D-74357 Boennigheim
          Germany
          Phone:
          Fax: +49-7143-22010
          Contact:
          EMail: cg@ssw.de
          WWW:

          Styx Technologies Inc.
          1132 Rhinecastle Way
          San Jose, CA 95120
          U.S.A.
          EMail: chris@styx.com

Last modified:

9.3.21 tgdb Description:

tgdb is a graphical user interface for gdb, the GNU debugger. Distribution Media:

Anonymous FTP, available as source package or Linux binary package. Licensing Policy:

Shareware OS Provisions:

Any Un*x-like OS (including Linux); requires X11R5. Documentation:

Comprehensive online/hypertext help (english); can be converted into various formats for printing. Product Support:

Updated when required; hotline and email support available. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

August, 1994. Countries with Distribution:

Worldwide (FTP). Price Range:

US$30 or DM50.00 per user; site licenses available. Installed Base: Vendor:

     HighTec EDV-Systeme GmbH
     Neue Bahnhofstr. 37
     D-66386 St. Ingbert, Germany
     Phone: +49 (0)6894/87 00 41
     Fax:   +49 (0)6894/87 00 44
     Email: tgdb@hightec.saarlink.de
     Contact: Michael Schumacher

Last modified:

September 9, 1994.

9.3.22 TowerEiffel

Description:

TowerEiffel is a complete software engineering tool for creating scalable systems in the object oriented programming language Eiffel. TowerEiffel for Linux includes a high performance Eiffel 3 compiler, open development environment, programming tools including debugger, browser and automatic documentation generation, and a base set of reusable software components. Key features include fast executable code, global system optimization, user controllable garbage collection, clear and precise error messages, exception handling, genericity, automatic system builds, automatic documentation generation and built in test support. A unique capability of TowerEiffel is Eiffel, C, and C++ interoperability.

Distribution Media:

4 Diskettes or QIC150 Tape

Licensing Policy:

TowerEiffel is licensed on a per user, right-to-use basis. There are no run-time fees. Two types of licenses are available:

1) Non-Commercial license for students and individuals developing software which will no be distributed for commercial use. Upon permission from Tower, software can be published as shareware.

2) Commercial license for commercial developers who will distribute or deploy developed software for commercial use.

OS Provisions:

no special requirements

Documentation:

500+ page user reference manual in English

Product Support:

60 days free technical support provided with purchase. Non-Commercial support is 30% per year of current retail price. Includes technical support via email or fax, and free upgrades (excluding shipping cost). Different support programs available for commercial developers.

Linux Support:

n/a Extra Features and Add-Ons:

TowerEiffel supports interoperability with C and C++. Motif GUI Libraries, and Data Structure Library (Booch) sold as add-ons.

Available Since:

November 1994 Countries with Distribution:

TowerEiffel is available worldwide with aurthorized distributors in UK, Ireland, Germany, Switzerland, Japan, and China

Price Range:

USD249 for non-commercial license; USD1295 for commercial license Installed Base:

N/A Vendor:

          Tower Technology Corporation
          1501 West Koenig Lane
          Austin, TX 78756, USA
          Phone:  800 285 5124 or +1 512 452 9455
          Fax:    +1 512 452 1721
          E-mail: tower@twr.com
          www: http://www.cm.cf.ac.uk/Tower/

Last modified:

April 5, 1995

9.4 Financial Software

All about financial tools.

9.4.1 BB Tool

Description:

BB Tool is a powerful stock charting, technical analysis and portfolio management tool. Distribution Media:

FTP download, or 3.5” diskettes Licensing Policy:

Per machine license. FREE demo version is available via FTP from ftp://ftp.portal.com/pub/ctor OS Provisions:

kernel version 1.1.52+, XFree86 3.1.1 Documentation:

PostScript formatted document is available via FTP download (file bb doc.ps.gz). The document has 42 pages including tutorial and reference guide. Man pages are also provided. BB Tool has extensive context sensitive online help. Available in English only.

Product Support:

Free patch of current version available via Email or FTP. Linux Support:

N/A Extra Features and Add-Ons:

Historic stock data on US Stock Exchange (NYSE, AMEX, NASDAQ) are available at USD1.00 per stock. Daily stock quote is available FREE, distributed via Email.

Available Since:

Sun SPARC version available since December 1993. Linux version available April 1995. Countries with Distribution:

USA Price Range:

Linux version: USD79; Sun SPARC version: USD89 Plus tax and shipping (if applicable) Installed Base:

N/A Vendor:

     Falkor Technologies
     P.O.Box 14201
     Fremont, CA 94539, U.S.A.
     Phone: +1 (510) 505-0700
     Contact: Henry Chen
     Email: ctor@shell.portal.com

Last modified:

May 10, 1995

9.5 Mathematics

Mathematics can as well be done with Linux ...

9.5.1 Maple V

Description:

Maple V Release 3 is a powerful general purpose computer algebra system. Maple V is a system for solving mathematical problems symbolically (instead of using paper and pencil and a lot of time) and numerically. Maple V’s graphics (various types of 2D, 3D, animation) visualizes the solutions. Maple V has also a programming language (Pascal like) which allows to extend the library of 2500+ functions.

Distribution Media:

3.5” disks Licensing Policy:

License for unlimited use in time of the bought version, floating license with user based pricing OS Provisions:

Linux kernel 1.0; XFree86 or Motif possible but not necessary; release number 2100: at least 8 MByte RAM suggested

Documentation:

3 books (”First Leaves: A Tutorial Introduction” 253 pgs., ”Language Reference Manual”, 267 pgs. and ”The Maple Handbook” 497 pgs.), Installation Manual. Manual pages of all functions and commands, help browser, keyword search. Documentation and help in English, some books in various languages exist.

Product Support:

updates/maintenance depending on license, please call the distributor or WMSI. Linux Support:

N/A Extra Features and Add-Ons:

share library via EMail and FTP. Free mailing list available (maple group@daisy.waterloo.edu). Available Since:

September, 1994. Countries with Distribution:

US/Canada, Europe, Asia Price Range:

German list price: DM 1785 (plus shipping and handling plus VAT). Please call the national distributor or WMSI.

Installed Base:

Vendor:

     Waterloo Maple Software
     450 Phillip Street, Waterloo, Ontario, Canada N2L 5J2
     Phone: +1 (519) 747-2373
     Fax:   +1 (519) 747-5284
     email: info@maplesoft.on.ca
     Germany, Switzerland and Austria:
     Scientific Computers GmbH
     Franzstr. 106
     D-52064 Aachen, Germany
     Phone: +49 (0)241/26041
     Fax:   +49 (0)241/44983
     E-Mail: info@scientific.de
     Contact: Monika Germ

Last modified:

Sep 19, 1994

9.5.2 Mathematica

Description:

Mathematica is a computational and visualization software package that’s used by engineers and scientists worldwide for quick, accurate numeric and symbolic answers.

Distribution Media:

3.5” 1.44MB DOS floppy disks

Licensing Policy:

N/A

OS Provisions:

Will take advantage of numeric coprocessor, if present. 20MB free disk space recommended (30MB required during installation). 8MB RAM required, 16MB RAM recommended for the Mathematica kernel.

Documentation:

Includes Mathematica : A System for Doing Mathematics by Computer, Installation Manual and User’s Guide for Unix computers, Guide to Standard Mathematica Packages, Release Notes, and additional technical reports. On-line help, on-line documentation.

Product Support:

Free technical support Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1995 Countries with Distribution:

U.S., Canada, Europe, Latin America, South Africa, Asia, Far East, Australia (everywhere) Price Range:

Starting at USD995 in the U.S. Installed Base:

N/A Vendor:

     Wolfram Research, Inc.
     Phone:  +1 (800) 441-6284 or +1 (217) 398-0700
     Fax:
     Contact:
     E-mail: info@wri.com
     WWW:    http://www.wri.com
     Europe:
     Wolfram Research Europe Ltd.
     Phone:  +44-(0)1993-883400
     E-mail: info-euro@wri.com

Last modified:

August 16, 1995

9.5.3 Matlab and Simulink

Description:

MATLAB is a high-performance, interactive numeric computation and visualization environment that combines the advantages of hundreds of packaged advanced math and graphics functions with a high-level language. SIMULINK is a powerful, interactive software package for modeling, analyzing, and simulating dynamic nonlinear systems.

Distribution Media:

CD-ROM Licensing Policy:

SiteID-Locked - Contact The MathWorks for details. OS Provisions:

Linux ’95 V1.2, 16 MB RAM, 20 MB disk space. Documentation:

Printed documentation, online help, MATLAB Reference Guide (HTML) Product Support:

The MathWorks supports a variety of Internet-based services to provide product information and services 24 hours a day. These services give immediate access to a wealth of resources, product data sheets, technical support, answers to frequently asked questions, user- contributed materials, open-forum discussions, and more. (For details ask vendor. hm)

Linux Support:

None. Extra Features and Add-Ons:

Toolboxes: Fuzzy Logic, Signal Processing, Image Processing, Spline, Symbolic Math, Neural Network, Nonlinear Control Design, Control System, Robust Control, Mu-Analysis & Synthesis, System Identification, Optimization, Statistics, Real-Time Workshop. Blocksets: Fixed-Point Blockset, DSP Blockset

Available Since:

June 1995 Countries with Distribution:

Spain, Portugal, United Kingdom, Belgium, The Netherlands, Luxembourg, Australia, Switzerland, Japan, Taiwan, Korea, Israel, Poland, Germany, France, Italy, Austria

Price Range:

Available upon request. Installed Base:

N/A Vendor:

     The MathWorks, Inc.
     24 Prime Park Way
     Natick, MA 01760, USA
     Phone:  +1 (508) 653-1415 x4322
     Fax:    +1 (508) 653-2997

     Contact: technical: Brian Bourgault, Marketing: Enza Froio
     E-mail: brian@mathworks.com, efroio@mathworks.com
     WWW:    http://www.mathworks.com

Last modified:

August 16, 1995

9.5.4 Reduce

Description:

REDUCE is a general purpose system for the symbolic manipulation of mathematical formulae (computer algebra) in science and engineering. More information via WWW from http://www.zib-berlin.de/Symbolik/reduce.

Distribution Media:

REDUCE for Linux is normally distributed on 3.5 inch diskettes. Tape cartridges of different formats are available on request.

Licensing Policy:

One time fee per installation. Discount for multiple licenses. Free demo version available via FTP from ftp://ftp.zib-berlin.de/pub/reduce/demo/linux.

OS Provisions:

Independent of LINUX kernel: the .c and .s sources for the LISP kernel are included and can be recompiled locally. X Windows not necessary but recommended. 8 Mbyte RAM recommended, better more.

Documentation:

A 230 page User’s manual is available. A reference manual is accessible in hypertext style on line under X Windows. All documents in English. The manual and additional documentation on disk in LaTeX.

Product Support:

Free patches of the current version available via electronic mail or ftp. Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1993 Countries with Distribution:

All countries except Iraq and Yugoslavia. Local distributors in Russia, Japan, Czec Republic, China.

Price Range:

Personal REDUCE: approx. USD99; Professional REDUCE: approx. USD499. Plus freight and tax. Installed Base:

several dozens Vendor:

          Konrad-Zuse-Zentrum Berlin
          Heilbronner Str. 10
          D-10711 Berlin-Wilmersdorf, Germany
          Phone:  +49 (0)30/89604-195
          Fax:    +49 (0)30/89604-125
          E-Mail: melenk@zib-berlin.de
          Contact: H.Melenk

Last modified:

Sep 28, 1994

9.6 Network Management

9.6.1 Galacticomm BBS

Description:

Worldgroup for UNIX, a client/server online software platform integrating workgroup/e-mail applications, commercial online services, and BBS technology.

Distribution Media:

CD-ROM, tape, 3.5” disks. Licensing Policy:

concurrent user-license, no free demo or shareware is available. OS Provisions:

1.1.59 and higher kernel, XFree86 3.1.x, 55MB disk usage, RAM depends on number of simultaneous users (approximately 750kb per user).

Documentation:

printed documentation, 550 pages, online client-side help. Product Support:

30-day free installation support, service contracts there- after. Classroom and/or on-site training available at extra costs.

Linux Support:

N/A Extra Features and Add-Ons:

too many to list.

Available Since:

Oct. 95 (previous versions of software since Aug. 94).

Countries with Distribution:

N/A

Price Range:

baseline 8-user system, USD1695.

Installed Base:

1000+

Vendor:

     Galacticomm, Inc.
     4101 SW 47 Avenue, Suite 101
     Ft. Lauderdale, FL 33314, USA
     Phone:  +1 (305) 583-5990, +1 (800) 328-1128
     Fax:    +1 (305) 583-7846
     E-mail: unix@unix.gcomm.com

Last modified:

Sept 14, 1995

9.6.2 NetEye

Description:

NetEye is a complete SNMP based Network Management System. NetEye’s main features are: Standard X-Window System and OSF/Motif user interface Autodiscovery and automapping of all IP based objects MIB Browser to simplify the navigation of all MIB-2 and private.enterprise variables Trouble ticketing system capable of storing and forwarding tickets via email and fax Capture, filtering and storing of SNMP device traps Topological, logical and spatial network views using color codes User definable bitmaps to represent network objects on maps User definable alarm and warning thresholds and time plots of any number of MIB variables Builtin address book to store and retrieve all fax and email addresses of other network managers Online hyper help for unexperienced operators

Distribution Media:

3.5” Floppy disk or ftp Licensing Policy:

Licensed per CPU, also available for OEM and source licensing. Full product is available on request for one month free of charge evaluation.

OS Provisions:

X11R5 and a video card capable of handling 256 colors. An ethernet card or PPP dialup connection. Minimum hardware requirements are 16 MByte RAM and a 386/40 CPU for managing small networks. Best performances using a 486DX2/66 CPU. Disk requirements: 12 MByte for basic installation that will grow depending on use and size of database of managed objects.

Documentation:

Documentation in english. Booklet of about 100 pages detailing all operations. Product Support:

email, fax, telephone. First 90 days free. USD450/year for ongoing support and updates. Linux Support:

N/A Extra Features and Add-Ons:

Comes with a compiled and ready to use copy of University Ingres 8.9 Available Since:

September 1993 Countries with Distribution:

UK, USA, Canada, Italy and worldwide (manufacturer) Price Range:

Complete network management system with trouble ticketing service: USD5500 list price. Source and OEM contracts: contact manufacturer.

Vendor:

     Soft*Star s.r.l.
     Via Camburzano 9
     10143 Torino, Italy
     Phone:  +39 11 746092
     Fax:    +39 11 746487
     Contact: Enrico Badella
     E-mail: softstar@pol88a.polito.it
             eb@relay1.iunet.it

Last modified:

November 1, 1994 9.6.3 Venus

Description:

Venus is an distributed administration tool for any network built of Unix workstations from different hardware vendors. A VENUS network presents itselfs to the user as a homogeneous computer system and garantees transparent access to any network resource. It provides software tools for major cluster management tasks: configuration management, software distribution, user and filesystem administration.

Distribution Media:

N/A Licensing Policy:

Commercial. OS Provisions:

Kernel 1.2, XFree86 3.1 Documentation:

Printed doumentation, online help. Product Support:

Hotline, email, courses ... on request. Linux Support:

Upon request. Extra Features and Add-Ons:

N/A Available Since:

1994 Countries with Distribution:

World. Price Range:

Upon Request. Special prices for educational institutions. Installed Base:

10 Linux Venus clients. (growing rapidly...) Vendor:

     science + computing GmbH
     Hagellocher Weg 71
     D-72070 Tuebingen, Germany
     Phone:  +49 (0) 7071/9457-0
     Fax:    +49 (0) 7071/9457-27
     E-mail: info@science-computing.uni-tuebingen.de
     WWW:    http://www.science-computing.uni-tuebingen.de/
     Contact: Olaf Flebbe

Last modified:

August 15, 1995

9.6.4 VU-BBS

Description:

Visually oriented BBS system and system/user administration package. VU BBS offers tools for installing and managing a large and active site as well as providing an easy to use and yet powerful commercial BBS environment.

Distribution Media:

Public distributions of VU BBS may be found at ftp.amaranth.com. Licensing Policy:

VU BBS is freely available at many major ftp sites. Our primary site is ftp.amaranth.com. In addition, annual registration is offered at USD45/year which provides immediate access to product updates and enhancements. VU BBS is distributed fully functional with no limitations on simulataneous user access, and is registered on a per-running-machine basis.

OS Provisions:

Linux kernel 1.1.47 or later. 4 megs of free space for installation. Running requirements are less than 200k per VU BBS user online!

Documentation:

Limited at this time. Product Support:

In addition to registration, a publically maintained mailing list for VU BBS sysops is hosted through majordomo@julie.pond.com.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

March 1995. Countries with Distribution:

world Price Range:

see above Installed Base:

N/A Vendor:

          Tycho Softworks
          25 Ceder St.
          Garfield, NJ 07026, USA
          E-mail: mycroft@mnsinc.com
          Contact: David Sugar

Last modified:

August 15, 1995

9.7 Text Processing

All kind of text processing, ASCII- and GUI-based

9.7.1 CRiSP

Description:

CRiSP is a graphical text editor on various Unix and Windows platforms which is 100% compatible with BRIEF. What makes Crisp different is that it tries to deliver its power in an intuitive point & click environment, without taking away the keyboard.

Distribution Media:

3.5” diskettes. Tapes available at extra charge. Licensing Policy:

Both node locked and floating licenses are available. Demonstration copy is archived at ftp://ftp.uu.net/vendor/vital.

OS Provisions:

RAM: 0.5 MByte, harddisk: 6 MByte (full), 2 MByte minimum. Openlook / Motif / Character versions are available on just about all commercial Unix platforms, Linux and Windows 3.1/NT.

Documentation:

printed, on-line manuals, on-line help, English Product Support:

updates and maintenance can be purchased at $100.00 US per license per year. It provides free upgrades for one year.

Linux Support:

We are currently giving away Slackware Professional 2.0 away for users ordering Crisp. This promotion however is only valid till end of October.

Extra Features and Add-Ons:

N/A Available Since:

March 1994 Countries with Distribution:

International Distributor: VITAL; Europe: Lasermoon Price Range:

USD99.99 (list: USD 149.99) without printed manuals. Printed user manual can be purchased at an additional USD40.00

Installed Base: Vendor:

     VITAL SOLUTIONS INC
     4109 Candlewyck Drive
     Plano, TX 75024, USA
     Phone:  +1 (214) 491-6907
     Fax:    +1 (214) 491-6909
     E-Mail: info@vital.com
     Contact: Gigi Mehrotra

Last modified:

Nov 1, 1995

9.7.2 ibgsXaed

Description:

ibgsXaed is an editor which is portable to every X Window System that offers OSF/Motif. The editor has normal edit functions and a lot of special features that support newcomers and professionals (f.e. function keys for system commands execution, variable fonts, undo, redo, online help, hotlist, iconbar, etc.).

Distribution Media:

3.5” diskette Licensing Policy:

free demo version OS Provisions:

X Window System X11R5, OSF/Motif 1.2, approx. 2 MByte harddisk Documentation:

printed manual, online help, both in English or German Product Support:

hotline, updates Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1994 Countries with Distribution: Price Range:

DM 799,- for one license, sliding prices for more licenses Installed Base: Vendor:

          ibgs GmbH
          Promenade 7
          D-52076 Aachen, Germany
          Phone:  +49 (0)2408/9455-53
          Fax:    +49 (0)2408/9455-15
          Contact: Mr.Sohn

Last modified:

Sep 28,1994

9.8 X Windows related products

Alternative X Windows servers, widget sets, window managers.

9.8.1 BXwidgets Description:

BXwidgets is a supplementary widget set to develop especially commercial Motif applications. It includes special widgets for tabulars, masks, formatted input field and the full support for a complete help system. All has Motif look and feel and is platform independent.

Distribution Media:

3,5” DOS diskette. Licensing Policy:

Source code licence. OS Provisions:

Any kernel, X11R5 or X11R6, Motif 1.1 or Motif 1.2. Documentation:

About 250 pages of documenation with examples. Several demo programs. Product Support:

Support and update service is available (1200,-DM or about 750,-US$ Linux Support:

N/A Extra Features and Add-Ons: Available Since:

1992. Countries with Distribution:

Germany, UK, USA. Price Range:

1600 DM (about 1000 US$ for a full source code licence. Installed Base:

N/A Vendor:

          BREDEX GmbH
          Fallersleber-Tor-Wall 23
          D - 38100 Braunschweig
          Germany
          Phone: +49 / 531 / 24 33 0 - 0
          Fax: +49 / 531 / 24 33 0 - 99

     Contact:
     EMail: info@bredex.de
     WWW: http://www.bredex.de

Last modified:

June 14, 1996.

9.8.2 BXwidgets/DB

Description:

BXwidgets/DB is a supplementary widget set to access relational databases. Data can be automatically queried form the database, displayed in various forms, changed and stored in the database. All applications built with BXwidgets/DB are database independent. The database access is performed by transaction-servers. Transaction-Server and database may reside on other hotsts in the network.

Distribution Media:

CD-ROM. Licensing Policy:

Per OS License, no run-time restrictions, Evaluation-License available. OS Provisions:

X11R5 and Motif1.2 must be installed. Documentation:

Printed manual in English. Product Support:

Support and update service is available. Linux Support:

N/A Extra Features and Add-Ons: Available Since:

1996. Countries with Distribution:

N/A Price Range:

DM 4900.

Installed Base:

N/A Vendor:

     BREDEX GmbH
     Fallersleber-Tor-Wall 23
     D - 38100 Braunschweig
     Germany
     Phone: +49 / 531 / 24 33 0 - 0
     Fax: +49 / 531 / 24 33 0 - 99
     Contact:
     EMail: info@bredex.de
     WWW: htpp://www.bredex.de

Last modified:

June 14, 1996.

9.8.3 Metrolink Motif

Description:

Motif 1.2.4 - Graphical User Interface - Complete Runtime and Development package with Color Pixmap Support.

Distribution Media:

3.5” diskettes (3), or FTP: The customer places the order with a credit card, then we email him our license agreement, installation instructions, release notes, and the FTP instructions – including the current password.

Licensing Policy:

Licensing is per CPU. NFS mounting not allowed. Site licenses are available for 10 or more copies. OS Provisions:

Linux 0.99pl15g or higher, 8-16 MByte RAM, 9 MByte HardDisk Documentation:

OSF User’s Manual (120 pages), man pages (electronic), English. Product Support:

90 days free tech support, 30-day money back guarantee, Extended Tech Support Contract $100 US per year.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

Motif 1.2.4 for Linux since August 1994 (Motif 1.2.2 for Linux since June 1993). Countries with Distribution:

USA, Germany, France. Price Range:

$149.00 US through October 31, 1994; $199 US after that if the special offer isn’t extended. Installed Base:

2,600 copies as of Sep 23, 1994 Vendor:

          Metro Link Incorporated
          4711 N. Powerline Rd.
          Fort Lauderdale, FL 33309, USA
          Phone:  +1 (305) 938-0283
          Fax:    +1 (305) 938-1982
          E-Mail: sales@metrolink.com (general info)
          Contact:  Holly Robinson (holly@metrolink.com, other
                  questions \& orders)

Last modified:

Sep 19, 1994

9.8.4 MoCheap Description:

A port of Motif to Linux, derived from and compatible with OSF/Motif 2.0. Requires Linux and X-windows and includes complete runtime and development system licensed for one CPU.

Distribution Media: Licensing Policy: OS Provisions: Documentation: Product Support: Linux Support: Extra Features and Add-Ons: Available Since: Countries with Distribution:

Price Range:

Installed Base:

Vendor:

     Yggdrasil Computing, Inc.
     4880 Stevens Creek Blvd.  #205
     San Jose, CA  95129-1034
     Phone: 408-261-6630
     Fax: 408-261-6631
     Contact:
     EMail: info@yggdrasil.com
     WWW:

Last modified:

9.8.5 OpenGL

Description:

OpenGL for Linux from Evans and Sutherland. OpenGL is a vendor neutral application programming interface developing 2D and 3D graphics applications that are highly portable to a variety of computing environments from PCs to workstations.

Distribution Media:

CD-ROM. Future updates will be made available via the World Wide Web.

Licensing Policy:

All of the OpenGL binaries are encrypted. A key for decryption is provided with the media kit. The encryption is not DES so shipment overseas is allowed.

OS Provisions:

OpenGL for Linux was been tested with kernel revisions from 1.0 to 1.2.8. It was compiled against XFree86-3.1.1 but you can also re-link your GLX server against the new XFree86-3.1.2. Recommended platforms: 486/33 ISA, 8 megabytes of ram, SVGA video. Best tested was a Pentium/90 PCI, 16 megabytes of RAM with an ATI Mach 64 PCI video.

Documentation:

Complete OpenGL man pages. References to OpenGL publications.

Product Support:

OpenGL for Linux from Evans and Sutherland is 79 U.S. dollars. Support is by e-mail only. This is to keep product cost down. The product passes conformance so very little problems have been reported. If you are currently running XFree86, you shouldn’t have any problems. As updates and enhancements or demos are released, they will be provided for free via WWW server.

Linux Support:

OpenGL for Linux isn’t bundled with Linux releases. Extra Features and Add-Ons:

It includes the complete XFree86-3.1.1 source and binary trees. It also includes a special OpenGL LinkKit. This LinkKit is used to reconfigure and re-compile your server if you need to. With this you can upgrade to new video drivers or add the GLX extension to XFree86-3.1.2 package.

Available Since:

OpenGL for Linux from Evans and Sutherland has been available since June of 1995. Countries with Distribution:

Even with encryption, shipments can be made to most overseas markets. Currently, the only distribution channel is Portable Graphics Inc. in Austin Texas.

Price Range:

The cost for this package is 79 U.S. dollars. Installed Base: Vendor:

     Portable Graphics
     P. O. Box 161002
     Austin, TX 78716
     Phone: 800-580-1160
     Phone: 1-512-306-0460
     Fax: 800-580-0616
     Fax: 1-512-306-0016
     E-Mail: linuxogl@portable.com

Last modified:

June 11, 1996.

9.8.6 X Inside X Servers

Description:

Replacement X Server for the XFree86 and other Servers for Linux. Supports Actix, ColorGraphics, Compaq, ELSA, Matrox, Number 9, ATI, Boca, Diamond, Orchid, Reveal, STB, TechWorks. Chipsets include Matrox, Number 9 Imagine-128, ATI Mach64, S3 964. Supports Mouse Systems, Microsoft, Logitech mice.

Distribution Media:

single floppy diskette Licensing Policy:

Per cpu license. Site licenses available.

OS Provisions:

Tested on versions of Linux between 1.0.8 and 1.1.18. No known problems with any Linux distribution. Requires around 3MB hard disk. Best used with a mouse...

Documentation:

Postscript formatted manual available via FTP from ftp://ftp.xinside.com/accelx/1.0/Manual. Hardcopy shipped with first copy of Server license.

Product Support:

Installation support included in price; available by email, Fax (and by phone if we have to). Update service available, US$100 for four revisions a year, to track new chipsets and other performance improvements. US$50 to upgrade to specified revision (e.g. 1.1 to 1.2). Product automatically sends email on failure. We may respond to a problem before you know you have it!

Linux Support:

N/A Extra Features and Add-Ons:

Various extensions will be offered, starting with PEX. Available Since:

May 94 Countries with Distribution:

Distribution currently in negotiation. Complete information available via FTP from ftp://ftp.xinside.com/accelx/1.0/All/prodinfo.txt. Europe: Delix Computers, Stuttgart

Price Range:

US$199.00. Student, academic and VAR discount programs. Student discount is 20% for Release 1.0 and 50% for Release 1.1. Post & Packing is between US$3 for US Mail to US$15 for Next Day delivery. Delix: Accelerated X DEM 159, Motif 2.0 DEM 259.

Installed Base: Vendor:

     X Inside Incorporated
     P.O. Box 10774
     Golden, CO 80401-0610, USA
     Phone:  +1 (303) 384-9999
     Fax:    +1 (303) 384-9778
     E-mail: info@xinside.com
     Europe:
     delix Computer GmbH

          Schloss-Str. 98
          D-70176 Stuttgart, Germany
          Phone: +49 711/621027-0
          Fax:   +49 711/613590
          E-mail: info@delix.de

Last modified:

April 7,1995

9.9 Other Software

Everything else that does not fit in one of the other sections by now.

9.9.1 Abuse

Description:

Abuse is a high-tech extremely fun action game using a classic side-scroller derivative format. Distribution Media:

The Registered Linux version is available on CDROM. The Shareware Linux version is available from http://crack.com

Licensing Policy:

The Shareware Linux version can be distributed without profit for free. The Shareware Linux version can be distributed for profit by asking Crack dot Com for permission. The Registered Linux version can be distributed for profit by opening a discussion with ddt@crack.com.

OS Provisions:

8Mb, X11R6, harddisk usage minimal. Documentation:

Online available with shareware and registered versions. Product Support:

E-mail support via abuse@crack.com. Phone tech support for registered versions only. Linux Support:

None. Extra Features and Add-Ons:

None. Available Since:

October, 1995.

Countries with Distribution:

Price Range:

US$45 + shipping from 1 800 810 0022. Can also order through abuse@crack.com. See shareware order.txt file for more details.

Installed Base:

Large but unknown.

Vendor:

     Crack dot Com
     PO Box 163025
     Austin, TX  78716-3025
     Phone: 1 800 810 0022
     Fax: 1 970 330 8773 (orders only)
     Contact: Dave Taylor
     E-mail: ddt@crack.com

Last modified:

June 13, 1996.

9.9.2 Amadeus music software

Description:

Fully professional music notation and printing software. Supports MIDI-interface for realtime and step input. Printer drivers to some matrix-, Ink-jet and Laserprinters. Supports postscript translation for output on PS-printers or use in Ghostscript for printing on any printer supported by GS. Supports TIFF (import and export). The functionality is somewhat roff or TEX look a like. ASCII-files (one per system) containing source are processed into files, containing one page each, wich can be converted into graphic files for viewing or editing on screen or printing, or into other formats as PS or TIFF. The software is highly automated in the actual note drawing. This makes it possible for blind musicians to print music for seeing if they use a blindsign-display to handle the computer.

Distribution Media:

Diskettes in tar-format (About 10)

Licensing Policy:

Strictly commercial. No demo at hand. Postcriptfile with detailed information and lots of music examples available. Post a message to m8770@abc.se and it will be sent to you.

OS Provisions:

Package contains kernel. XFree86 needed (not for blind users). 8Mb RAM, about 20 Mb HD.

Documentation:

Doc. in English and German.

Product Support:

Support via Tel. or Fax. Support free if the information not is available in the manuals. Uppdates about once a year, prices varies.

Linux Support:

Only if it is strictly related to the software

Extra Features and Add-Ons:

Are developed if requested. Price related to the work needed.

Available Since:

March 1994

Countries with Distribution:

Directly to the intire world from the developer except for Scandivia. For Scandinavian customers: see Vendor

Price Range:

DEM 4000; Blind version DEM 6500; Upgrade from the Idris-version (Atari) DEM 2000

Installed Base:

N/A

Vendor:

          Wolfgang Hamann
          Amadeus Notenzats
          Winterstr. 5
          D-81543 Muenchen, Germany
          Phone:  +49 89 669678
          Fax:    +49 89 669579
          Contact: Wolfgang Hamann
          Scandinavian customers and e-mail:
          Jerker Elsgard
          MIDIBIT AB
          Box 161
          S-618 23 KOLMARDEN, Sweden
          Phone/Fax: +46 11 391663
          Contact: Jerker Elsgard
          E-mail:  m8770@abc.se

Last modified:

April 5, 1995 9.9.3 aqua zis

Description:

(Time Series Information System). A system to store and maintain huge amounts of measurement data (mainly hydrological data). Time series with several 100,000’s of values can be handled. You may retrieve, edit and insert data very fast. A lot of statistical and arithmetical operators are applicable. Time series may be visualized in a full scale GUI, includeing high performance zooming ans scrolling. The user may define individual GUI’s in an easy script language and with the time series programming language Azur. Reports on any paper size including full graphics are possible through Azur. Output is produced in Postscript and HPGL formats.

Distribution Media:

We install at your site. Demo disk available (tar) Licensing Policy:

machine or site licenses available. Test licenses are limited in time. OS Provisions:

Linux version 0.99pl6 or later. Motif required if you use the GUI features. When set up as a time series database server, no graphics is needed. RAM usage 8 – 16 MB. Disk space required: some 15 MB for executables and help files. For the data: several GB’s.

Documentation:

printed manuals covering end user and admin tools (English version soon to appear). Program documentation and design available only in English.

Product Support:

12 months warranty, maintenance contract recommended. One day training at your site included. Linux Support:

N/A Extra Features and Add-Ons:

aqua trop precipitaion reconstrucion tool using GIS methods. DigiTul digitizing of precipitaion, water-level etc. paper registrations. aqua log gound water data management.

Available Since:

Jan 1994 Countries with Distribution:

Germany, Austria, soon EC. Price Range:

machine license: 39,000 DM, multiple machines: graduated discounts, site licenses on request, reasonable discout for universities.

Installed Base:

some 4 Linux-only systems are running. Vendor:

     aqua_plan
     Ing.-Ges. fuer Problemloesungen
     in Hydrologie und Umweltschutz mbH
     Mozartstr. 16
     D-52064 Aachen, Germany
     Phone:  +49 (0)241/31430
     Fax:    +49 (0)241/31499

Last modified:

Aug 26, 1994

9.9.4 ChemApp

Description:

ChemApp is an object library of subroutines that enables a user to apply the powerful calculation capabilities of ChemSage via an easily programmable interface. It fills the need for a flexible tool to allow ChemSage complex chemical equilibrium calculations for multicomponent, multiphase mixtures to be included into other programs. ChemApp can be used to design application-specific programs for process simulation, modeling or control; or to develop your own teaching program. It can be added as a module to third-party process simulation packages; for example, to commercial CFD programs, to general simulation programs, or to your own company’s program. For more information, visit our World Wide Web site.

Distribution Media:

1 diskette. Licensing Policy:

ChemApp is available for general use under licence, and under royalty terms for integration into third-party packages.

OS Provisions:

Kernel version >= 1.2.13 (might work with earlier versions too) Documentation:

Printed documentation (English). Product Support:

Full product support (incl. update service, technical support via EMail, fax, phone). Prices for maintenance, training, and thermochemical data for use with ChemApp on request.

Linux Support:

Only for problems concerning Linux ChemApp. Extra Features and Add-Ons:

Prices for custom additions to ChemApp (e.g. user-specific solution models) upon request. Available Since:

Spring 1996. Countries with Distribution:

Worldwide, manufacturer in Germany, distribution partners in USA, Finland, Japan, France, U.K., Czech Republic.

Price Range:

Annual licence: DM 5,000 - 11,000 Perpetual licence: DM 12,500 - 27,500 Reduced prices for universities and recognized national research institutions.

Installed Base:

N/A Vendor:

     GTT-Technologies
     Kaiserstrasse 100
     52134 Herzogenrath
     Germany
     Phone: +49-2403-59533
     Fax: +49-2403-59661
     Contact: Bob Fullerton-Batten (licencing),
             Dr. Stephan Petersen (technical)
     EMail:
     WWW: http://gttserv.lth.rwth-aachen.de/~sp/tt/

Last modified:

June 24, 1996.

9.9.5 CLiX

Description:

CliX is a OSF/Motif based desktop available for almost any UNIX platform, providing support for remote and local file and directory manipulations, handling print and batch queues and message based communication.

Distribution Media:

Free Demo without network and batch functionality available via FTP from ftp://ftp.science-computing.uni-tuebingen.de/pub/clix-light/. Floppy version on request.

Licensing Policy:

Free demo without remote functionality, full commercial version available on request. OS Provisions:

Kernel 1.2, XFree86 3.1, 16 MB RAM min Documentation:

Full version: printed documentation, approx 150 pages, online help. Free version: german docu without figures on ftp server.

Product Support:

No support for free version. Commercial version: Hotline, classes.. on request Linux Support:

Not appropriate Extra Features and Add-Ons:

Not appropriate Available Since:

May 1995. Countries with Distribution:

World. Price Range:

Free Demo. Full Version (Linux) DEM 1000. Domain licenses available. Special discount (approx. 70 %) for educational sites.

Installed Base:

N/A Vendor:

          science + computing GmbH
          Hagellocher Weg 71
          D-72070 Tuebingen, Germany
          Phone:  +49 (0) 7071/9457-0
          Fax:    +49 (0) 7071/9457-27
          E-mail: info@science-computing.uni-tuebingen.de
          WWW:    http://www.science-computing.uni-tuebingen.de/
          Contact: Olaf Flebbe

Last modified:

August 15,1995 9.9.6 Emulus Description:

Emulus is an X-Windows/Motif application which uses TCP/IP to establish a connection to an IBM mainframe host, emulating a 3270 terminal. Emulus supports mainframe graphics, including 3270 graphics datastream extension used by mainframe SAS Software. Other features include customization through dialogs, a scripting facility to record and playback host interaction, EasyPads (customized windows of buttons), and cut and paste between Emulus and other X applications. Emulus also comes with Helplus, an hypertext help viewer modeled after WinHelp. It is used to provide online help for Emulus, but can also be used to develop Help for other X applications.

Distribution Media:

CD-ROM only. Licensing Policy:

Licensing is per CPU. OS Provisions:

Linux 1.2 or higher, 8-16 MByte RAM, 5 MByte HardDisk; Motif statically linked. Documentation:

Printed book, man pages, on-line help, English. Product Support:

Technical Support from SAS Institute. Linux Support:

N/A Extra Features and Add-Ons: Available Since: Countries with Distribution: Price Range:

US $99; includes hardcopy documentation, and Technical Support for one year. Installed Base:

N/A Vendor:

          SAS Institute Inc., USA
          Book Sales Division
          Phone: (919) 677-8000
          Fax:
          Contact: Phil Herold
          E-mail: saspjh@unx.sas.com
          WWW: http://www.sas.com

Last modified:

June 13, 1996.

9.9.7 Executor

Description:

Commercial Macintosh Emulator. Fast CPU Emulation (75 MHz 486DX4 approximates 25 MHz 68040). Allows many Macintosh applications, including Word 5 and Excel 4 to run without requiring anything from Apple. Reads, Writes and Formats 1.4 MB Macintosh formatted floppies. Reads and Writes SCSI Macintosh disks. Reads Macintosh formatted CD-ROMs. Color support, Prints to PostScript printers. Has demo applications so you can run it with no Macintosh experience.

Distribution Media:

Electronic now, CD-ROM when 2.0 ships

Licensing Policy:

Executor for Linux can be found in ftp://ftp.ardi.com/pub/ExecutorLinux. It has a ten minute time limit that can be removed via a serial number and authorization key. There are enough incompatibilities that potential customers are strongly encouraged to examine the Executor in its time limited demo form before ordering. However, we offer a 30 day unconditional money back guarantee.

OS Provisions:

Is known to work with kernel 1.2.0. Standard distribution takes about 7 MB. Needs 8 MB of its own RAM to run quickly. The demo version is dynamically linked with the XFree86 3.1 libraries and with libc/libm-4.5.26.

Documentation:

English FAQ at ftp://ftp.ardi.com/pub/executor-faq.ascii.

Product Support:

Free upgrades to licensees of 1.99<x>, up to and including the first official release of Executor for Linux: Executor/Linux 2.0.

Linux Support:

questions@ardi.com

Extra Features and Add-Ons:

N/A

Available Since:

Has been in pre-beta since March of 1995. Is expected to be officially released in November of 1995.

Countries with Distribution:

Current pre-beta releases are available only directly from us. Within six months of 2.0’s release we will probably have distribution worldwide.

Price Range:

Pre-beta copies are available for USD49 for students, faculty and staff of educational institutions, USD99 for everyone else. These prices are good at least until September 15. Somtime soon thereafter the suggested retail price for the non-educational license will be USD249. The educational price will rise significantly, also.

Installed Base:

129 Linux users as of August 12, 1995. Many more DOS and NEXTSTEP users.

Vendor:

     ARDI
     Suite 4-101
     1650 University Blvd., NE
     Albuquerque, NM 87102, USA
     Phone: +1 505 766 9115
     Fax: +1 505 247 1899
     Contact:
     E-mail: questions@ardi.com, orders@ardi.com, bugs@ardi.com

Last modified:

August 15, 1995.

9.9.8 Mazama Packet Filter

Description:

The Mazama Packet Filter is a Linux-based IP packet filtering or firewall application.

Distribution Media:

The Mazama Packet Filter package is distributed on 1 (one) floppy (3.5”) disk. Also included is the latest version of the InfoMagic CD-ROM Linux distribution.

Licensing Policy:

Evaluation copies are available on request, and will be distributed based solely on our evaluation of the potential client. Source code will not be distributed with the eval version (but with the production version. hm)

OS Provisions:

kernel version: 1.2.1 through 1.2.8, support for later versions available as they surface. We track kernel versions closely. RAM: min. 8MB, suggested 16+MB. Hard disk: MPF software < 2MB. Recommend total at least 250MB, (accounts for log file disk space requirements). The MPF package includes the InfoMagic Linux distribution which includes all necessary source for the kernel. The MPF installation also provides patches for versions 1.2.0 through 1.2.8, (see above) to update the kernel for compatibility with the MPF.

Documentation:

Printed and bound documentation, as well as online html version. Online version available at Mazama Web Site <http://www.mazama.com/>. English language only currently.

Product Support:

Free updates via anonymous ftp. E-mail and phone support. Training and consulting available on request. Inquire about prices.

Linux Support:

Linux support, training, and configuration assistance available on request. Once again, inquire about prices.

Extra Features and Add-Ons:

none. Available Since:

May 15, 1995 Countries with Distribution:

We are currently negotiating with several foreign distributors. Price Range:

Mazama Packet Filter: USD750. Mazama Enterprize Firewall: USD1500 Installed Base:

5 copies currently in circulation. Vendor:

     Mazama Software Labs, Inc.
     15600 NE 8th St., Suite B1 #544
     Bellevue, WA 98008
     Phone: +1 (206) 545-1808
     Fax:
     Contact: Eric D. Berg
     E-mail: info@mazama.com (general inquiries),
             support@mazama.com (tech support)

Last modified:

June 9, 1995 9.9.9 PROCHEM-C

Description:

PROCHEM-C is an integrated Software System for users dealing with design, construction, maintenance, operation and control of plants in the fields of

- Pipeline Construction

- Chemical & Pharmaceutical Industry

- Power Plant Construction

- Petrochemical Industry

- Food Processing

- Environment Technology Distribution Media:

DAT or QIC150 Licensing Policy:

per Workstation or Server-Licence OS Provisions:

kernel version 1.0 , modified portmapper/mount XFree86 Version 2.1, Motif Version 1.2.3, RAM : 32 - 64 MByte (min. 16 MByte) DISK: 1 GByte (min. 500 MByte) Swap 120 MByte (min. 80 MByte)

Documentation:

ON-Line integrated in PROCHEM-C (printed optional) ca. 1700 Pages. German or English. Product Support:

Update Service : All 3 Months or on User-request. Prices for Maintenance and Training on request. Linux Support:

Only for Problems concerning LINUX <–> PROCHEM-C Extra Features and Add-Ons:

N/A Available Since:

1983, LINUX Version since 1993 Countries with Distribution:

N/A Price Range:

DEM 2,000 to DEM 65,000 (i.a.w. amount of features. hm) Installed Base:

N/A Vendor:

     COMPLANSOFT CAD GmbH
     Sulzbacher Strasse 15 - 21
     D-65812 Bad Soden, Germany
     Phone:  +49 6196/56 06-0
     Fax :   +49 6196/56 06-66
     Contact : Chris Chirila

Last modified:

April 6, 1995

9.9.10 ROSIN NC

Description:

ROSIN-NC-Werkbank (ger: workbench) The ROSIN-NC-Werkbank is a comprehensive set of tools which accomplish a CAD system. These tools comprise the ROSIN-VDAFS-Processor to process digitised data, and data in the VDAFS format; the ROSIN-Post Processors to generate 2- to 5-axis milling programs for all major CNC process controls; the ROSIN-NC-Editor to analyse and edit existing NC programs; the ROSIN-NC-Visualiser to visualise the tool path and generate printed information about a NC program; the ROSIN-Engraver to mill logos or fonts of any kind.

Distribution Media:

Tapes (DAT, QuarterInch) or floppy disk

Licensing Policy:

All tools of the ROSIN-NC-Werkbank are independently licensed via the ROSIN License System, the customer gets a production license for each purchased tool. The combination of license types with different tools is possible. There are individual licenses for single users, 2-5 users etc. Please consult vendor for more information.

OS Provisions:

Kernel Version higher or equal to 0.99pl14; Motif; RAM 16 MByte, recommended 20 MByte; the necessary disk space for a complete installation with all tools is 25 MByte. It is recommended to ensure a free disk space of at least 1,5 times the size of the input file(s).

Documentation:

An Online Help Function is available with all tools. With the purchase of each tool a manual is supplied in the language chosen by the customer (languages available: German, English, French and Italian, other languages on request). Size of manuals: Depending on the tool from 60 to 170 pages each.

Product Support:

Maintenance contracts are offered for all tools and cost 1,3% per month of the license price. With a maintenance contract the customer gets regular product updates and manual updates. Migration to a different hardware platform is significantly cheaper (3% of the license price instead of 12% for customers without maintenance contract). The telephone hotline is free of charge. Our telephone hotline is available from 8.30 to 17.00 MET, languages are German and English. We offer special prices for product installation and training courses (75% of the normal service prices) for product customers. Prices for product updates can be provided on request.

Linux Support:

(there was no ”official” statement by ROSIN, however, I was told on the phone that ROSIN ships complete systems with pre-installed software, and that they would offer limited Linux support as far as handling their software is concerned. hm)

Extra Features and Add-Ons:

N/A

Available Since:

1994

Countries with Distribution:

Germany, France and Mexico (please consult vendor for more information). We are looking for distributors in the UK, Spain and Italy.

Price Range:

The lower prices (in DM) are for Single User Licenses, the higher ones for Multi User Licenses, floating type. Prices for all other licenses are included in this price range. ROSIN-VDAFS-Processor 7500 to 9750 ROSIN-3-axis Post Processor 4500 to 11050 ROSIN-5-axis Post Processor 13500 to 32500 ROSIN-NC-Editor 7500 to 9750 ROSIN-NC-Visualiser 3500 to 4550 ROSIN-Engraver 1500 to 1950

Installed Base:

Vendor:

     ROSIN Gesellschaft fuer technische Datenverarbeitung mbH
     Roniger Weg 13
     D-53545 Linz am Rhein, Germany
     Phone:  +49 (0)2644/97003-0
     Fax:    +49 (0)2644/97003-32
     Contact: Gisela Buechner (Marketing)

Last modified:

Sep 19, 1994 9.9.11 SPATCH Description:

SPATCH software allows alphanumeric paging from a Linux system. With SPATCH, Linux users can send text messages to alphanumeric pagers from a user, an application, an operating system, or an E-Mail System.

Distribution Media:

3.5” Disk, File Transfer Licensing Policy:

30 Day Evaluation Kits are available. OS Provisions:

N/A Documentation:

SPATCH includes printed documentation for installation and use. Product Support:

Installation support is provided free of charge. Linux Support:

N/A Extra Features and Add-Ons:

SPATCH SECURITY (USD149) Used by ISPs to prevent unauthorized access to SPATCH. SPATCH CALL COUNT REPORTING (USD99) Used by ISPs for reporting and billing of call counts.

Available Since:

July 1995 Countries with Distribution: Price Range:

USD199 - USD999 Installed Base:

Currently SPATCH is installed on 2 Linux systems. Vendor:

          The Hyde Company, Inc.
          P.O. Box 900190
          Atlanta, GA 30329, USA
          Phone:  +1 (770) 495-0718
          Fax:    N/A
          Contact: Alan Lewis
          E-mail: spatch@cy.com

Last modified:

August 15, 1995

9.9.12 Unix Cockpit

Description:

The Unix Cockpit (UC) is a nifty new file manager for Unix/X11 that smoothly integrates a directory tree, file browsers, custom menus and the classic Unix shells into one highly customizable productivity tool of a kind long missed on Unix systems.

Distribution Media:

anon. ftp from ftp.uu.net (192.48.96.9), directory /vendor/UniX11/linux. Licensing Policy:

On Linux, USD25 shareware. For the people who wrote Linux, free. OS Provisions:

Kernel 1.0.8 or higher. Small, single file executable (500k). No installation necessary. Documentation:

Fully indexed, postscript version of tutorial is available via ftp. Tutorial is also available online from within UC. UC has full context sensitive online help. The language used is English.

Product Support:

Small updates and bug fixes are free and made available via ftp. Extra support is via email courtesy support hotline at support@UniX11.com. Special training/installation help on request.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

The Linux version is beta testing since 11/94. The SGI versions have been shipping since 9/94. SunOS, HP & DEC Alpha versions are in beta, too.

Countries with Distribution:

All distribution goes over the internet. Price Range:

USD25 shareware. Installed Base:

It is still a beta version, which is free, so no Linux copies have sold. Quite a few beta testers, though.

Vendor:

     Henrik Klagges
     UniX11 Software Development Corp
     Moorbachweg 7
     D-83209 Prien, Germany
     E-mail: henrik@UniX11.com

Last modified:

April 5, 1995

9.9.13 Unix Auto-Tutor

Description:

Unix Auto-Tutor is an automatic Unix tutor – a comprehensive, fully interactive, self-teaching CBT system for Unix training. It is organized like a traditional textbook with hypertext pages. Currently it has 9 chapters and 76 sections with 87 interactive tutorial sessions attached to its hypertext pages. More than 150 Unix commands, such as find, grep, sort, sed, awk and vi, are introduced. Can be used to provide in-house training or on-line training through BBS or Internet.

Distribution Media:

3.5” diskettes. Licensing Policy:

Per CPU with the number of users purchased, eg. 2 user edition would limit 2 users at a time. OS Provisions:

Kernel version .99 or higher, runable under xterm, Motif version is under construction, 300k RAM and 1.2M harddisk usage.

Documentation:

A 185+ page hardcopy of the on-line hypertext, online help available, all in English. Product Support:

90 days free telephone and e-mail support. Instructor-led Unix training based on the Auto-Tutor is also available in house.

Linux Support:

No support for Linux OS itself. Extra Features and Add-Ons:

Additional copies of the textbooks are available at USD25 each. Available Since:

March 1995 Countries with Distribution:

Direct sales world wide.

Price Range:

Single User Edition: USD199 per CPU (for individual only). USD299 per CPU (for business). Add USD100 for each additional user. Shipping is additional.

Installed Base:

N/A

Vendor:

     Sunny Micro Lab
     1300 Britannia Road East, Suite #208
     Mississauga, Ontario L4W 1C8, Canada
     Phone: +1 (905) 795-9292
     Fax:   +1 (905) 795-9291
     E-mail: info@sunni.com  or  sunni@hookup.net
     Contact:  Mo Chen

Last modified:

August 15, 1995

9.10 Free Software for Commercial Hardware

This section is somewhat “exotic”. It covers free Linux software (drivers et cetera) for specialized hardware. This is no “commercial software”, however, the software is unusable without the hardware.

9.10.1 Cyclades Cyclom-Y

Description:

Cyclom-Y is serial multiport card for ISA/EISA PC’s. 8 and 16-port models are available. There are RJ-12 and DB-25 connector options. It is based on the Cirrus Logic CD1400 serial processor. This RISC-based processor supports on-board character processing and allows speeds up to 115.2 kbps. Drivers for Xenix, SCO, ISC and UnixWare are available from Cyclades. There is a driver for Linux available in the Internet. Board variants: Cyclom-8Ys: 8 ports, RJ-12 connectors (no RTS signal); Cyclom-8Yo: 8 ports, DB-25 connectors in a octopus cable (full-modem); Cyclom-8Yb: 8 ports, DB-25 connectors in a external box (full-modem); Cyclom-8Yb+: 8 ports, same as 8Yb with surge protection; Cyclom-16Y: 16 ports, DB-25 connectors in a external box (full-modem).

Distribution Media:

3.5” diskettes

Licensing Policy:

The Linux device driver software is free and is available via FTP from ftp://sunsite.unc.edu/pub/Linux/kernel/patches/misc.

OS Provisions:

(the driver is available as a patch relative to Linux 1.1.8. A driver for newer kernels is under development. hm)

Documentation:

The user’s manual contains all the reference information to configure and install the card. Specific information about the Linux driver can be found in readme files.

Product Support:

SW updates from Internet. Cyclades will provide technical support on the hardware, but the driver is a freeware and is provided ”as is”. Cyclades makes no warranties about the suitability of this software for any purposes. Some companies bundle the driver in their Linux distributions and can help in the installation and configuration of the card.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

July 1994 Countries with Distribution:

worldwide Price Range:

     List Price (end user):
     CYCLOM 8YS \$459
     CYCLOM-8YO \$549
     CYCLOM-8YB \$599
     CYCLOM-8YB+ \$699
     CYCLOM-16Y/DB25 \$794
     Promotion for resellers only (first-time buyer):
     CYCLOM 8YS \$99
     CYCLOM-8YO \$199
     CYCLOM-16Y/DB25 \$399
     NOTE:  All prices are F.O.B. Fremont, CA, USA

Installed Base: Vendor:

     Cyclades Corporation.
     44140 Old Warm Springs Blvd
     Fremont, CA 94538, USA
     Phone:  +1 (510) 770-9727 or (800) 347-6601
     Fax:    +1 (510) 770-0355
     E-Mail: cyclades@netcom.com
     Contact: Doris Li (Sales and Marketing).
     Distributor for Germany:
     roka EDV und Datenkommunikationsberatung GmbH
     Elbestr. 25
     D-47800 Krefeld, Germany
     Phone: +49 (21 51) 49 75-10
     Fax:   +49 (21 51) 49 75-32
     Contact: Bernward Averwald
     E-Mail: bernward@roka.net

Last modified:

Sep 24, 1994

9.10.2 Stallion Technologies Multiport Serial Boards

Description:

(Stallion Technologies sent me a vast amount of data concerning their 4 thru 64 port boards. Too much to print here, really. Basically, there’s the EasyIO series with 4 or 8 ports, the EasyConnection 8/32 with 8 to 32 ports and EasyConnection 8/64 with up to 64 connections. The Stallion Linux drivers also support the older ranges of Stallion Technologies boards, including the ONboard, Brumby and original Stallion familys of boards. Please contact vendor for more information. hm)

Distribution Media:

3.5” diskettes for supported operating systems (including SCO UNIX and XENIX, UnixWare, SVR4.x, Interactive Unix, NetWare Connect V1.0 or higher, NACS V3.0 and higher, MS-DOS, OS2). Linux drivers available from Internet archive sites, stallion.com, sunsite.unc.edu, tsx-11.mit.edu and their mirrors. (Note the Linux driver is not supported by Stallion Technologies).

Licensing Policy:

The Linux driver is available at no charge and is copyrighted under the GNU General Public License.

OS Provisions:

The most current driver supports kernel versions 1.2.5 and later. The drivers are updated regularly to keep pace with kernel changes.

Documentation:

Hardware reference documentation provided with the boards. The The Linux driver package contains a README file with driver installation and usage information.

Product Support:

Stallion Technologies does not support the Stallion board Linux drivers. They were developed independently of Stallion Tecnologies. Stallion Technologies will provide support for their hardware.

Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

October 1994 Countries with Distribution:

Worldwide distribution. Available in most countries. Price Range:

Prices for the serial boards vary from country to country. Contact your nearest Stallion Technolgies office for local availability and pricing. The driver package contains a file with all Stallion Technologies offices.

Installed Base:

N/A Vendor:

          Stallion Technologies Pty Ltd.
          56 Sylvan Rd
          Toowong, QLD, 4066, Australia
          Phone: +61 7 270 4242
          Fax:   +61 7 371 8881
          E-mail: allsales@stallion.com

Last modified:

June 9, 1995

9.11 German corner

This section contains software which is only marketed in Germany and/or other German-speaking countries or which is otherwise specific for Germany or German-speaking countries. The companies listed here sent me entries in German... Dieser Abschnitt enthaelt Software, die nur in Deutschland und oder anderen deutschsprachigen Laendern vertrieben wird, oder die anderweitig spezifisch fuer den deutschen oder deutschsprachigen Raum ist. Die hier gelisteten Firmen schickten mir deutsche Eintraege...

9.11.1 BBSD-Druckerei

Description:

BBSD-Druckerei bietet alle erdenkliche Funktionalitaet, die zur Verwaltung in kleinen und mittleren Druckereien benoetigt wird:

• Verwaltung von Papieren zur Kalkulation und zur Lagerverwaltung

• Kundenverwaltung

• Erstellung von Kalkulationsboegen

• Erstellung von Angeboten mit Staffelpreisen etc.

• Erstellung von Auftragstaschen

• Erstellung von Rechnungen

• Drucken von Etiketten

• Listen ueber alle erfassten und verwalteten Daten

BBSD-Druckerei arbeitet mit einer Textoberflaeche, so dass fuer die meisten Arbeitsplaetze nur einfache ASCII-Terminals angeschafft werden muessen. Allerdings existieren auch Schnittstellen zu PC’s, so dass eine problemlose Datenuebernahme - bspw. aus der Kundenverwaltung in eine PC-Textverarbeitung (Serienbriefe!?) - moeglich ist.

Distribution Media:

Disketten, DAT Licensing Policy:

Die SW wird nur kommerziell vertrieben. Eine Praesentation kann nach Absprache bei uns oder bei dem Kunden vorgenommen werden.

OS Provisions:

Kernel Version 1.2, mind. 8 MB RAM und 0,5 MB RAM je Benutzer Documentation:

Deutsches Handbuch, Online-Hilfe Product Support:

Update-Service, Schulung, persoenliche Betreuung, Fernwartung Linux Support:

N/A Extra Features and Add-Ons:

N/A Available Since:

Juni 1995 Countries with Distribution:

Deutschland Price Range:

DEM 2000 (abhaengig von kundenspezifischen Anpassungen) Installed Base:

(am 02.08.95) ¿= 4 Vendor:

  Baastrup \& Bayard GbR Software Design
  Mallinckrodtstr. 7
  44145 Dortmund
  Phone:  +49 (0)231/981 20-60
  Fax:    +49 (0)231/981 20-62
  E-Mail: bbsd@bbsd.de

Last modified:

September 14, 1995

9.11.2 Orgaplus

Description:

Orgaplus vertreibt eine ganze Auswahl branchenspezifischer Loesungen:

• J5 - Projektmanagement & Controlling fuer Agenturen und Projektierende Unternehmen (ab DM 6000,-)

• netConnect - Kabelmanagement & Endstellenverwaltung (ab DM 5500,-)

• R4 - Warenwirtschaft fuer den Reifenhandel (ab DM 6500,-)

• W5 - Warenwirtschaft (ab DM 6500,-)

• S2 - Technik / Wartung / Service (ab DM 3500,-)

• M3 - Menue- & Security-Management entspr. ISO 9000 Audit Vorschriften (ab DM 850,-)

• F4 - Finanzbuchhaltung (ab DM 1750,-)

• APAS - Lohn & Gehalt mit Voll-DUEVO. AOK-Systemgeprueft (ab DM 2800,-)

• A7 - Vertrieb- & Marketingmanagement (ab DM 1250,-)

• T5 - Textverarbeitung (integriert in die opus Software) (ab DM 300,-)

• G4 - Reportgenerator fuer den Sachbearbeiter (ab DM 1500,-)

• V2 - Versandpapierschreibung, Zollabwicklung (ab DM 4500,-)

• B2 - Anlagenbuchhaltung (ab DM 2500,-)

• K4 - Kostenrechnung (ab DM 4500,-)

• E2 - Elektronische Archivierung (ab DM 20000,-)

• Data Dictionary (ab DM 4500,-)

• Klassenbibliothek (ab DM 8500,-)

• Programmgenerator (ab DM 12500,-)

• Maskengenerator Text + GUI (ab DM 2500,-) Distribution Media:

Disketten, DAT Licensing Policy:

Demonstration erfolgt in unserem Haus oder nach Vereinbarung, Kurzpraesentation ueber WWW in Erstellung.

OS Provisions:

Kernel Version 1.2, XFree86 alle Versionen, Motif nicht noetig, mind. 8 MByte RAM, mind. 30 MByte freien Speicherplatz.

Documentation:

Deutsches Handbuch, selbst erweiterbare Online-Hilfe Product Support:

Update-Service, Schulung, persoenliche und telefonische Betreuung, Fernwartung, Fernschulung. Preise siehe beiliegendem Honorarverzeichnis.

Linux Support:

Installation von Linux, komplette Einrichtung von Linux. Preisgestaltung nach Aufwand, wegen unterschiedlichen Kundenanforderungen.

Extra Features and Add-Ons: Available Since:

Maerz 1995 Countries with Distribution:

Deutschland, Oesterreich, Schweiz, Niederlande Price Range:

User- und modulabhaengige Staffelung siehe Preisliste. Installed Base: Vendor:

          orgaplus Software GmbH
          Tel.:   +49 71 31/59 75-0
          Fax:    +49 71 31/59 75-99
          E-mail: info@orgaplus.de
          WWW:    http://www.orgaplus.de
          Contact: Herr Michael Walz (Geschaeftsfuehrer),
                   Herr Jaeger (Vertrieb)

Last modified:

10. Mai 1995

10 Support

Who you’re gonna call ...

This section contains some companies who provide for Linux support. The raw format is as follows:

Company name
Address
Phone
Fax
EMail
Contact
Type of support (via phone, e-mail, inhouse, training etc.)
sample prices

10.1 Companies

The following list is virtually alphabetically sorted.

• ARIS Technology, Inc.

Mississippi Research & Technology Park, Suite 204 Starkville, MS 39759, USA Phone: +1 (615) 325-2319 Fax: +1 (615)323-3726 EMail: info@aris.com Contact: Dr. Michael L. Stokes Type of support: Local Area networking support, Linux/UNIX/Windows compatibility software, Custom software, Web Authoring, Internet Access(Broadband 65kpbs and above) Sample prices: Technical support (USD25/hour). For Internet prices, see http://www.aris.com

• Baastrup & Bayard GbR Software Design Mallinckrodtstr. 7 D-44145 Dortmund, Germany Phone: +49 (0)231/981 20-60 Fax: +49 (0)231/981 20-62 E-Mail: bbsd@bbsd.de Contact: Joern Baastrup or Bjoern Bayard Type of support: We offer phone/EMail/inhouse support, installation, ongoing administration, training, consulting and support for porting existing applications to LINUX. Sample prices:

Installation/Administration support (via phone/EMail) DM 100,00 / hour Developer support DM 135,00 / hour Training (contracts with better conditions are possible) DM 125,00 / hour Consulting DM 140,00 / hour Specialities: Development of business applications, Workflow-Managment based on COSA by Software Ley GmbH, WWW and EMail, ”Plug & Play” LINUX-Workstations, Cyclades Multiportcards, ISDN with LINUX

• Bitbybit Information Systems Radex Complex, Kluyverweg 2A NL-2629 HT Delft, The Netherlands Phone: ++31-15-682569 Fax: ++31-15-682530 EMail: bitbybit@runner.knoware.nl Contact: Ir. N. Simon Type of support: On-site, Network Sample prices: Hfl 120,- per hour. Specialities: Installation, Internet Connectivity. Consulting, Custom-Made Software, Timetabling Systems, Database Interoperability.

• ClearVu TWS 95 Clearview Drive Christiansburg, VA 24073, USA Phone: +1 (703) 231-3938 EMail: Thomas Dunbar Sample prices: EMail/phone USD70/hr; inhouse USD70/hr + traveling expenses Speciality: custom typesetting systems

• Couvares Consulting 146 Mill Lane Amherst, MA 01002, USA Phone: +1 (413) 253-2589 EMail: couvares@family.hampshire.edu Contact: Peter F. Couvares Type of support: We offer phone/EMail support, installation, ongoing administration, training, and specialized consulting for free software and Linux (and other UNIX) systems. Sample prices: USD40/hour commercial, USA20/hour nonprofit, sliding scale for personal

• Gregg Weber 1076 Carol Lane #3 Lafayette, CA 94549, USA Phone: +1 (510) 283-6264 EMail: gregg@netcom.com Contact: Gregg Weber Type of support: ? Sample prices: USD40/hour, or per job quote. Free initial discussion of job requirements.

Specialities: Installation, problem solving, C and assembly language programming, Embedded controller microcoding, cumputer systems and peripheral electronic design and development. Systems integration. DOS, MS-Windows, X Window System, Novell, Lantastic, internet EMail

• Frontier Internet Services Limited Allied House 45 Hatton Garden London EC1N 8EX, England phone +44 (0)171 242 3383 fax +44 (0)171 242 3384 EMail: info@ftech.net Contact: Gareth Bult, Piers Cawley Type of support: All. Sample prices: UKP250/day Specialities: Internet Services Providers Systems/Network Integraters (Unix/Novell/Windows) General Consultancy

• Le Reseau netwerksystemen BV Burg. F. v. Ankenweg 5 NL-9991 AM Middelstum The Netherlands Phone: +31 595 552431 Fax: +31 595 551279 EMail: support@reseau.nl WWW: http://www.reseau.nl Contact: Arthur Donkers Support: Phone, EMail, InHouse, software porting, software development, plug-and-play Linux machines Price: Support subscriptions for phone and EMail costs Dfl. 495/year, InHouse support costs Dfl. 75 / hour. Projects are done on a fixed price basis Extra: Le Reseau also sells Linux CD Roms by InfoMagic, Red Hat, Walnut Creek and Yggdrasil. Le Reseau is an Caldera Business Partner for the Netherlands. Le Reseau is also the home for the Dutch linux mailinglists

• LEMIS Schellnhausen 2 D-36325 Feldatal, Germany Phone: +49 (0)66 37/91 91 23 Fax: +49 (0)66 37/91 91 22 EMail: lemis@lemis.de Contact: Greg Lehey Type of support: Phone, EMail, inhouse Sample prices Hotline: DEM 3 per minute or part thereof; Contract: DEM 150/hour

• Michael Koehne Bremer Strasse 117 D-28816 Brinkum, Germany Phone: +49 (0)421 / 87 55 00 Fax: +49 (0)421 / 87 55 51 EMail: kraehe@bakunin.north.de (private) Type of support: N/A Sample prices: to be negotiated Specialities: Linux installation for database or internetworking use. OS/2 and WfW conectivity. Applications for Getraenkegrosshaendler, Buchhandlungen, Wasserwirtschaftsaemter.

• Randysoft Software 286 Cresta Vista Way San Jose, CA 95119, USA Phone: +1 (408) 229-0119 EMail: randy@randysoft.com Contact: Randy Hootman Type of support: Via phone, e-mail, inhouse, training, software development. C/C++, Perl, Shell, networking, internet, installation. Sample prices: Prices vary according to type of work and length of contract.

• Signum Support AB Box 2044 S-580 02 Linkoping, Sweden Phone: +46 (0)13 21 46 00 Fax: +46 (0)13 21 47 00 EMail: info@signum.se Contact: Magnus Redin, Inge Wallin Type of support: Phone, fax and EMail support Sample prices: Linux on CDROM with manual and installation support for one system, SEK 1,380; Internet server with WWW, Gopher, FTP, Sendmail, News/INN and DNS with administrationing tools that fits a pc system administrator and 40h support, SEK 30,000 Specialities: Linux, GCC, CVS version management, Emacs, Tcl/Tk and Internet services: WWW, Gopher, FTP, Sendmail, News/INN

• The Silk Road Group, Ltd. 13628 Bent Tree Circle #201 Centreville, VA 22020, USA Phone: +1 (703)222-4243 Fax: +1 (703)222-7320 EMail: bass@silkroad.com Contact:Tim Bass Sample Prices: subject to project and availability Specialities: IP network management specialists using Linux platforms for Internet Information Services, including httpd servers, DNS and news servers, SMTP mail relay hosts, IP firewalls. Large scale system and network design and integration for large corporations and goverment agencies. Clients include the USAF and the National Science Foundation. More info on http://www.silkroad.com/ • Systems Software

22 Irvington Cres Willowdale, On, Canada M2N 2Z1 Phone: +1 (416) 225-1592 EMail: Drew Sullivan Sample prices: $90 Canadian/hour Specialities: Unix Systems & Internet Consulting

• Sytek s.r.l.

Via Cernaia 14 I-10122 Torino, Italy Phone: +39 11 9884111 Fax: +39 11 9951487 EMail: mario@sytek.it, info@sytek.it Contact: Mariotti Mario Type of support: by e-mail, inhouse/outhouse, training Sample prices approx. USD60 per hour Specialities: General Unix Consultant, X11/Motif Experts, we own a GUI Builder for Motif (Xad), we have great experience in writing Unix Device Drivers (char & block).

• Wizvax Communications 1508 Tibbits Ave Troy, NY 12180 USA Phone: +1 (518) 273-4325 Fax: +1 (518) 271-6289 Contacts: Richard Shetron (owner/manager), Stephanie Gilgut (Linux/SunOS/Unix wizard) Support: We prefer to do as much support as possible via e-mail, phone, telnet. Sample prices: Prices vary depending on both level of service desired (24 hr/7 day, business hours, phone, EMail, etc.) and term of contract (one-shot, monthly, yearly, etc.). Typical prices range from USD45/hr for classes for beginners taught by JR level staff to USD120/hr for one-shot emergency calls to fix systems broken by incorrectly installed software/hardware by non-customers with no support contracts. Specialities: We primarily support the multiport serial boards and cables we sell. Support for products we sell is free for installation and initial setup. We can do Linux consulting, training, seminars, etc.